### v2

Abstract
Across a wide array of disciplines, many researchers use machine learning (ML) algorithms to identify a subgroup of individuals who are likely to benefit from a treatment the most (“exceptional responders”) or those who are harmed by it. A common approach to this subgroup identification problem consists of two steps. First, researchers estimate the conditional average treatment effect (CATE) using an ML algorithm. Next, they use the estimated CATE to select those individuals who are predicted to be most affected by the treatment, either positively or negatively. Unfortunately, CATE estimates are often biased and noisy. In addition, utilizing the same data to both identify a subgroup and estimate its group average treatment effect results in a multiple testing problem. To address these challenges, we develop uniform confidence bands for estimation of the group average treatment effect sorted by generic ML algorithm (GATES). Using these uniform confidence bands, researchers can identify, with a statistical guarantee, a subgroup whose GATES exceeds a certain effect size, regardless of how this effect size is chosen. The validity of the proposed methodology depends solely on randomization of treatment and random sampling of units. Importantly, our method does not require modeling assumptions and avoids a computationally intensive resampling procedure. A simulation study shows that the proposed uniform confidence bands are reasonably informative and have an appropriate empirical coverage even when the sample size is as small as 100. We analyze a clinical trial of late-stage prostate cancer and find a relatively large proportion of exceptional responders.
Key Words: causal inference, exceptional responders, heterogeneous treatment effects, treatment prioritization, uniform confidence bands
1 Introduction
In a diverse set of fields, machine learning (ML) algorithms are frequently used to identify a subgroup of individuals who are likely to benefit from a treatment the most (so called “exceptional responders”) or those who are negatively affected by it. In clinical studies, for example, identifying those who experience surprisingly long-lasting positive health outcomes when a majority of patients do not, is useful for understanding why a treatment work and determining which patients should receive it (e.g., Takebe et al., 2015; Conley et al., 2020). Similarly, it is critical to identify a subpopulation of patients who are harmed by a commonly used treatment. In public policy, finding those individuals who are likely to be helped by a treatment is a key methodological step for improving targeting strategies (e.g., Imai and Ratkovic, 2013; Chernozhukov et al., 2019).
An increasingly popular approach to this subgroup identification problem consists of two steps. Researchers first employ an ML algorithm to estimate the conditional average treatment effect (CATE) or its proxy (e.g., biomarker, baseline risk) given a set of pre-treatment covariates. Next, they use the resulting treatment prioritization score to select those individuals who are predicted to benefit from or harmed by the treatment. For example, researchers sort individuals based on their estimated score and then identify, as exceptional responders, a group of top scoring individuals who exhibit the largest positive group average treatment effect. We follow the literature and refer to the average treatment effect for group sorted by a treatment prioritization score as the GATES (Chernozhukov et al., 2019)
As a motivating study, we consider a well-known randomized clinical trial described by Byar and Green (1980) concerning the effect of administering a type of synthetic estrogen, called diethylstilbestrol, to prolong total survival among late-stage prostate cancer patients. The trial was overall unsuccessful as the treatment group with the highest dosage, 5.0mg, registered an average of 0.3 month decrease in total survival compared to the placebo group. However, given that the biological pathway of estrogen in reducing prostate cancer growth is well-known, clinical researchers believed that at least a certain subgroup of patients could have benefited from the treatment. Therefore, many researchers have attempted to find such a subgroup that exhibits a large GATES, by first applying ML, optimization, or heuristic methods to create a treatment prioritization score and then optimize over the threshold of the score with a certain objective function (e.g. Bonetti and Gelber, 2000; Rosenkranz, 2016; Ballarini et al., 2018; Bertsimas et al., 2019).
Despite its intuitive appeal and frequent use in practice, this approach faces several methodological challenges. First, there exists a considerable degree of statistical uncertainty in estimating the GATES for the selected subgroup. The treatment prioritization scores are often based on noisy and imperfect estimates of the CATE or other target quantities. Second, the subgroup of exceptional responders (or those who are negatively affected by the treatment) is usually small, and to make the matter worse, the size of experimental data used to identify such individuals is often not large either. Third, accounting for this statistical uncertainty is difficult because utilizing the same data to both evaluate the GATES and select the group of potential exceptional responders suffers from a multiple testing problem. Although this problem can be alleviated by setting aside a large holdout set for evaluation, such an approach is often not a feasible solution given the limited sample size of randomized control trials.
In this paper, we develop uniform confidence bands (intervals) for estimating the GATES among a subgroup of individuals defined by a quantile of treatment prioritization score. This group represents the subgroup of individuals who are predicted to benefit most from (or be harmed most by) the treatment. The proposed uniform confidence bands are valid regardless of the way in which the quantile value is chosen, thereby providing a statistical guarantee for a generic objective function to be optimized when selecting the group of individuals who are either most positively or most negatively affected by the treatment. For example, our methodology enables researchers to choose the quantile cutoff such that the GATES among the selected exceptional responders exceeds a certain threshold with a pre-specified probability.
The proposed methodology is based on Neyman’s repeated sampling framework (Neyman, 1923), assuming only randomization of treatment and random sampling of individual units without requiring any modeling assumption or repeated sampling methods. Thus, our methodology is applicable to any generic treatment prioritization score, including the ones that may not have any performance guarantee. Although we do not pursue in this paper, additional assumptions, such as monotonicity of treatment effect in treatment prioritization score, can be imposed to further obtain tighter confidence bands at the expense of generality (e.g., Chernozhukov et al., 2010). Our simulation study shows that the proposed uniform confidence bands are reasonably informative and provide an adequate empirical coverage for sample sizes that are as small as $n=100$ (Section 3).
To develop our methodology (Section 2), we first show that the problem of finding uniform confidence bands is equivalent to deriving the distribution of a particular functional of induced order statistics. This functional is also called a concomitant of order statistics in the literature, and has been the subject of independent interest (e.g., David and Galambos, 1974; Yang, 1977; Sen, 1976). In particular, Bhattacharya (1974) and Davydov and Egorov (2000) explore the asymptotic properties of general functionals of induced order statistics.
While the resulting asymptotic distributions are in general difficult to characterize, we show that the GATES are non-negatively correlated with one another regardless of how the treatment prioritization score is constructed. This allows us to uniformly bound the GATES by a Wiener process through an application of Donsker’s theorem. Although the resulting uniform confidence bands are not unique, we propose a specific choice that yields a narrow band for a reasonable range of quantile values. We further extend our theoretical results beyond the GATES to other causal quantities of interest, widening the applicability of the proposed methodology.
Finally, we apply the proposed methodology to the aforementioned randomized clinical trial (Section 4). We find that, with 95% of confidence, modern ML algorithms are able to identify a subset of patients whose survival would be increased by the treatment up to 27 months on average. This contrasts with a previous study which found, albeit without a statistical guarantee, a group of exceptional responders whose survival is increased by only up to 18 months on average (e.g., Bertsimas et al., 2019). This empirical application demonstrates that our methodology can exploit the power of modern ML algorithms while providing a statistical performance guarantee.
Related Literature.
The proposed methodology is related to several strands of the existing literature in computer science, optimization, and statistics. First, there exists a large literature whose goal is the development of methods to identify subgroups with large treatment effects using ML algorithms (e.g. Kehl and Ulm, 2006; Su et al., 2009; Foster et al., 2011; Hardin et al., 2013; Bertsimas et al., 2019; Cho and Ghosh, 2021). However, these existing works do not provide any statistical performance guarantee.
Second, Chernozhukov et al. (2019) considers the estimation of the GATES based on the data from randomized experiments. Although they use resampling methods to provide statistical inference, the resulting confidence bands are point-wise and are not uniformly valid across the quantile values. While Imai and Li (2022) improves upon this methodology by using Neyman’s repeated sampling framework (thereby, avoiding computationally intensive resampling methods), the authors also only provide point-wise confidence bands. Our work fills this methodological gap by developing uniform inference for the GATES estimation.
Finally, there exist a small but fast growing body of literature that develops statistical methods for experimentally evaluating the performance of data-driven individualized treatment rules (ITRs). Imai and Li (2023) use Neyman’s repeated sampling framework to show how to experimentally evaluate the efficacy of ITRs. Building on the up-lift modeling literature (e.g., Radcliffe, 2007), the authors propose the Population Average Prescriptive Effect (PAPE) curve as a performance measure of an ITR relative to the non-individualized treatment rule that treats the same proportion of randomly selected individuals. This is a causal analogue of the Receiver Operating Characteristic curve. Like this paper, Imai and Li (2023) provide an assumption-free statistical inference methodology to evaluate the performance of ITRs. Unlike our proposed methodology, however, their confidence bands are point-wise and do not yield uniform inference for the entire curve.
Similarly, Yadlowsky et al. (2021) considers the performance evaluation of ITRs, using a related quantity called the Targeted Operating Characteristic (TOC) curve. The TOC curve is similar to the prescriptive effect curve but compares the performance of an ITR with the treatment rule that treats everyone. While the authors show how to quantify statistical uncertainty for the area under the TOC curve, their inferential methods are point-wise and are not uniformly valid for the entire TOC curve. Unlike these two existing works, the proposed methodology yields uniform statistical inference, enabling the selection of exceptional responders and other subgroups with uniform statistical guarantees. As shown below, the proposed uniform confidence bands can be applied to the PAPE and TOC curves as well.
2 The Proposed Methodology
2.1 Setup and Assumptions
Suppose that we have a total of $n$ units independently sampled from a (super)population $\mathcal{P}$. We consider a classical randomized experiment where a binary treatment $T_{i}\in\{0,1\}$ is completely randomized with a total of $n_{1}$ units assigned to the treatment condition and the remaining $n_{0}=n-n_{1}$ units assigned to the control condition. For each unit $i=1,\ldots,n$, we observe the outcome $Y_{i}\in\mathcal{Y}$ as well as a vector of pre-treatment covariates, $\mathbf{X}_{i}\in\mathcal{X}$. The observed outcome for unit $i$ is given by $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$ where $Y_{i}(t)$ represents the potential outcome under the treatment condition $T_{i}=t$. This implies the stable unit treatment value assumption or SUTVA (Rubin, 1990). In particular, the outcome of one unit is assumed to be not affected by the treatment condition of other units. We summarize our setup here:
Assumption 1 (Random Sampling of Units)
We assume independently and identically distributed observations,
$\{\mathbf{X}_{i},Y_{i}(1),Y_{i}(0)\}_{i=1}^{n}\ \stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}\ \mathcal{P}$
Assumption 2 (Random Assignment of Treatment)
We assume complete randomization of treatment assignment,
$\Pr(T_{i}=1\mid\{\mathbf{X}_{i},Y_{i^{\prime}}(1),Y_{i^{\prime}}(0)\}_{i=1}^{n})=\frac{n_{1}}{n}\quad\text{for all }i=1,\ldots,n,\ \text{and }n_{1}=\sum_{i=1}^{n}T_{i}$
Without the loss of generality, we assume that a positive treatment effect is desirable. We consider a scoring rule $s(\cdot)$ that determines treatment prioritization where a greater score represents a higher priority. Formally, a scoring rule maps the pre-treatment covariates to a scalar variable,
$f:\mathcal{X}\longrightarrow\mathcal{S}\subset\mathbb{R}.$ (1)
The most commonly used scoring rule is the estimated conditional average treatment effect (CATE) given the pre-treatment covariates, i.e., $f(\mathbf{x})=\widehat{\mathbb{E}}(Y_{i}(1)-Y_{i}(0)\mid\mathbf{X}_{i}=\mathbf{x})$, using a machine learning (ML) algorithm. Another common scoring rule is the risk score, which is typically represented by the baseline conditional outcome, i.e., $f(\mathbf{x})=\widehat{\mathbb{E}}(Y_{i}(0)\mid\mathbf{X}_{i}=\mathbf{x})$.
Throughout, we assume that a scoring rule is given. This means that we do not account for the uncertainty that may arise from the estimation of scoring rule. However, we remain agnostic about how the scoring function is constructed. This means that the scoring rule may be a heuristic rule or derived from complex ML algorithms. In other words, the scoring rule may or may not accurately sort units based on the degree to which they benefit from the treatment. The only assumption we impose is that the resulting treatment prioritization score is continuous.
Assumption 3 (Continuous Treatment Prioritization Score)
The cumulative distribution of the treatment prioritization score $F(s)=\Pr(S_{i}\leq s)$ is continuous.
In practice, we can break a tie, if it exists, by adding a small amount of noise to the score though our framework does not account for the uncertainty that arises from this random tie-breaking procedure.
Next, we make two regularity assumptions. First, we assume that the CATE, i.e., $\mathbb{E}(\psi_{i}\mid\mathbf{X}_{i})$ where $\psi_{i}=Y_{i}(1)-Y_{i}(0)$, is continuous in the quantile of the treatment prioritization score:
Assumption 4 (Continuity of the conditional average treatment effect)
The conditional average treatment effect given the quantile of the treatment prioritization score $\mathbb{E}[\psi_{i}\mid F(S_{i})=p]$ is continuous for $p\in[0,1]$.
A similar assumption is commonly invoked in the causal inference literature (e.g. Wager and Athey, 2018; Yadlowsky et al., 2021). In addition, we assume that the individual treatment effect has a finite second moment. The assumption is necessary for our asymptotic convergence result to have a well defined distribution.
Assumption 5 (Moment Condition for the Individual Treatment Effect)
We assume that the individual treatment effect $\psi_{i}$ has a finite second moment, i.e., $\mathbb{E}(\psi_{i}^{2})<\infty$.
2.2 Subgroup Identification with a Statistical Performance Guarantee
We consider the setting where a researcher uses the above treatment prioritization score to identify a subpopulation of individuals who benefit from treatment the most. The same methodology can be applied (by simply flippting the treatment prioritization score via $S_{I}=-f(\mathbf{X}_{i})$) to the case where researchers wish to identify a subgroup of individuals who are harmed most by treatment.
Formally, we wish to find the quantile $p^{\ast}$ of the treatment prioritization score such that the sorted group average treatment effect (GATES) is maximized,
$p^{\ast}=\operatornamewithlimits{argmax}_{p\in[0,1]}\Psi(p)\quad\text{where}\ \Psi(p)\ =\ \mathbb{E}[\psi_{i}\mid F(S_{i})\geq p],$ (2)
where, under Assumption 2, the GATES is identified as,
$\Psi(p)=\mathbb{E}[Y_{i}\mid T_{i}=1,F(f(\mathbf{X}_{i}))\geq p]-\mathbb{E}[Y_{i}\mid T_{i}=0,F(S_{i})\geq p].$ (3)
If $\Psi(p)$ is known exactly, the optimization problem in Equation (2) leads to the selection of individuals, based on the quantiles of prioritization score, whose average treatment effect is the greatest. In practice, however, we only have a data set of finite sample size and $\Psi(p)$ may be estimated with bias and noise.
Specifically, researchers must solve the following empirical version of this optimization problem,
$\hat{p}_{n}\ =\ \operatornamewithlimits{argmax}_{p\in[0,1]}\widehat{\Psi}_{n}(p),$ (4)
where
$\widehat{\Psi}_{n}(p)=\frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]}\quad\text{and }\ \hat{\psi}_{[n,i]}=\frac{T_{[n,i]}Y_{[n,i]}}{n_{1}/n}-\frac{(1-T_{[n,i]})Y_{[n,i]}}{n_{0}/n}$ (5)
with the subscript $[n,i]$ denoting the $i$th order statistic where the ordering is based on the treatment prioritization score $S_{i}=f(\mathbf{X}_{i})$. Thus, for example, if we use $S_{[n,i]}$ to denote the $i$th order statistic of $S$, then we have $S_{[n,1]}\geq S_{[n,2]},\ldots,\geq S_{[n,n]}$.
While we can easily compute $\hat{p}_{n}$, the resulting estimate does not take into account for the estimation uncertainty of $\widehat{\Psi}_{n}$. This is problematic especially when the sample size is limited and/or the true proportion of exceptional responders $p^{\ast}$ is small. The problem is exacerbated by the fact that in practice the estimated CATE is often biased and noisy. In addition, because the selection of the parameter $p$ depends on the same data, the standard confidence band for $\widehat{\Psi}_{n}$ computed by conditioning on $\hat{p}_{n}$ suffers from a multiple testing problem and does not have a proper coverage guarantee.
To address these challenges, we consider the identification of exceptional responders with a statistical performance guarantee. We first develop an asymptotic uniform one-sided confidence band $C_{n}(p,\alpha)$ that covers the true value $\Psi(p)$ for any value of $p$ with the probability of at least $1-\alpha$, i.e.,
$\mathbb{P}\left(\forall p\in[0,1],\ \Psi(p)\geq\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha)\right)\geq 1-\alpha.$ (6)
We then use this uniform confidence band to obtain a statistical guarantee when selecting those who are predicted to benefit most from the treatment. Specifically, instead of selecting the value of $p$ that maximizes the estimated GATES as done in Equation (4), we choose the value of $p$ such that the lower confidence bound of the estimated GATES is maximized,
$\underline{\hat{p}}_{n}\ =\ \operatornamewithlimits{argmax}_{p\in[0,1]}\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha).$ (7)
This procedure gives a statistical guarantee that, with probability at least $1-\alpha$, the true GATES among those who are identified as benefiting most from the treatment is at least as great as the optimized lower confidence bound,
$\mathbb{P}\left(\Psi(p^{\ast})\geq\widehat{\Psi}_{n}(\underline{\hat{p}}_{n})-C_{n}(\underline{\hat{p}}_{n},\alpha)\right)\geq\mathbb{P}\left(\Psi(\underline{\hat{p}}_{n})\geq\widehat{\Psi}_{n}(\underline{\hat{p}}_{n})-C_{n}(\underline{\hat{p}}_{n},\alpha)\right)\geq 1-\alpha.$
Since this statistical guarantee does not depend on how the value of $p$ is selected, we can accommodate a wide variety of ways in which this threshold parameter is selected. For example, one may wish to identify the largest group of individuals whose GATES is at least as great as the pre-specified threshold $c$, i.e.,
$p^{\ast}(c)\ =\ \sup\{p\in[0,1]:\Psi(p)\geq c\}$
Suppose that we choose the value of $p$ such that the lower confidence bound matches with the threshold $c$, i.e.,
$\underline{\hat{p}}_{n}(c)\ =\ \sup\{p\in[0,1]:\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha)\geq c\}.$
Then, this procedure provides a statistical guarantee that, with the probability at least $1-\alpha$, the GATES for the selected group of individuals is at least as great as the threshold $c$.
$\mathbb{P}\left(\Psi(p^{\ast}(c))\geq\widehat{\Psi}_{n}(\underline{\hat{p}}_{n}(c))-C_{n}(\underline{\hat{p}}_{n},\alpha)\geq c\right)\geq 1-\alpha.$
Next, we show how to construct the uniform confidence band given in Equation (6).
2.3 Constructing Uniform Confidence Bands
We begin by bounding the finite sample bias and deriving the exact variance of the GATES estimator $\widehat{\Phi}_{n}$ given in Equation (5). Define $\widehat{F}_{n}(s)$ as the empirical CDF of the treatment prioritization score $S_{i}=f(\mathbf{X}_{i})$ based on the sample of size $n$. Furthermore, let $\hat{c}_{n}(p)\ =\ \inf\{c\in\mathbb{R}:\sum_{i=1}^{n}\mathbf{1}\{S_{i}>c\}\leq np\}$ denote the estimated $p$-quantile. Then, the application of Theorem 1 of Imai and Li (2023) gives the following results about the pointwise bias and variance of the GATES estimator in the current context.
Proposition 1 (Bias Bound and Exact Variance of the GATES Estimator)
Under Assumptions 1 and 2, the bias of the proposed estimator given in Equation (5) can be bounded as follows,
$\mathbb{P}(|\mathbb{E}\{\widehat{\Psi}_{n}(p)-\Psi(p)\mid F(\hat{c}_{n}(p))\}|\geq\epsilon)$
$\leq$ $1-B\left(p+\gamma_{p}(\epsilon),np,n-np+1\right)+B\left(p-\gamma_{p}(\epsilon),np,n-np+1\right)$
for any given constant $\epsilon>0$ where $B(\epsilon,\alpha,\beta)$ is the incomplete beta function (if $\alpha\leq 0$ and $\beta>0$, we set $B(\epsilon,\alpha,\beta):=H(\epsilon)$ for all $\epsilon$ where $H(\epsilon)$ is the Heaviside step function), and
$\gamma_{p}(\epsilon)\ =\ \frac{\epsilon p}{\max_{c\in[p-\epsilon,\ p+\epsilon]}\mathbb{E}(\hat{\psi}_{i}\mid F(S_{i})=p)}.$
The variance of the estimator is given by,
$\mathbb{V}(\widehat{\Psi}_{n}(p))\ =\ \frac{1}{p^{2}}\left\{\frac{\mathbb{E}(S_{p1}^{2})}{n_{1}}+\frac{\mathbb{E}(S_{p0}^{2})}{n_{0}}-\frac{p(1-p)}{n-1}\mathbb{E}(\hat{\psi}_{i}\mid\widehat{F}(S_{i})\geq p)^{2}\right\},$
where for $t\in\{0,1\}$ we have $S_{pt}^{2}=\sum_{i=1}^{n}(\hat{\phi}_{ip}(t)-\overline{\hat{\phi}_{p}(t)})^{2}/(n-1)$ with $\hat{\phi}_{ip}(t)=\mathbf{1}\{\widehat{F}(S_{i})\geq p\}Y_{i}(t)$, and $\overline{\hat{\phi}_{p}(t)}=\sum_{i=1}^{n}\hat{\phi}_{ip}(t)/n$.
To develop the uniform convergence result, we begin by showing that a normalized version of $p\widehat{\Psi}_{n}(p)$ converges in distribution to a zero-mean Gaussian process. In our proof, we utilize a generalized version of Donsker’s invariance principle for correlated random variables. For completeness, we state this theorem in Appendix Theorem S1, and refer readers to Borovkov and Sahanenko (1983) for its proof in the independent case and Ossiander (1987) for the correlated case. We now state the convergence result.
Proposition 2 (Convergence of the Normalized GATES Estimator)
Consider the random continuous polygon $s_{n}$ connected by the following basic points $b_{i}$ for $i=1,\cdots,n$:
$b_{i}=\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right).$
Under Assumptions 15, $s_{n}$ converges in distribution to a zero-mean Gaussian process uniformly in the Skorokhod space.
Proof is given in Appendix S1.
To develop the uniform convergence result, we consider the correlation between $\widehat{\Psi}_{n}(p)$ and $\widehat{\Psi}_{n}(q)$ for any $p>q$. After normalizing them by their group size, we can write the difference between these two GATES estimators as,
$p\widehat{\Psi}_{n}(p)-q\widehat{\Psi}_{n}(q)\ =\ \frac{1}{n}\sum_{i=\lfloor nq\rfloor}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]}.$
where $\hat{\psi}_{[n,i]}$ is defined in Equation (5). Therefore, the correlation between $\widehat{\Psi}_{n}(p)$ and $\widehat{\Psi}_{n}(q)$ is determined by the correlation among $\hat{\psi}_{[n,i]}$. If the treatment prioritization score $S_{i}=f(\mathbf{X}_{i})$ exactly equals the CATE, then $\psi_{[n,i]}$ will be strongly positively correlated with one another because they are a noisy version of the true CATE.
Somewhat surprisingly, however, regardless of the accuracy of this score, one can show that the sorted individual treatment effect estimates $\hat{\psi}_{[n,i]}$ are non-negatively correlated with one another.
Proposition 3
(Non-negative Correlation of Sorted Individual Treatment Effect Estimates) Under Assumptions 15 and for all $1\leq i<j\leq n$, we have:
$\operatorname{Cov}(\hat{\psi}_{[n,i]},\hat{\psi}_{[n,j]})\geq 0.$
Proof is given in Appendix S2. The above inequality is tight and reaches its bound of 0 when the score $S_{i}$ is independent of the individual treatment effect $\psi_{i}$, i.e., when the ML estimate of the CATE is completely uninformative. We emphasize that this inequality is general and applicable to arbitrary evaluation metrics beyond GATES. This non-negative correlation is induced by the nature of the ordering operation.
Figure 1: Simulation study illustrating a non-negative correlation between the treatment prioritization score $S_{i}$ and the individual treatment effect $\psi_{i}$. Even when the score is negatively correlated with the individual treatment effect, the individual treatment effects ordered by the score are positively correlated. The magnitude of positive correlation increases as the relationship between the score and individual treatment effect becomes stronger. The correlation becomes zero when the score is independent of the individual treatment effect.
To illustrate this property, we simulate the data using $n=100$ and $Y_{i}(t)=X_{i}(1+t)$ for $t=0,1$, where $X_{i}\sim N(0,1)$, implying $\psi_{i}=X_{i}$. Then, we create the treatment prioritization score using $S_{i}=rX_{i}+\sqrt{1-r^{2}}Z_{i}$ where $Z_{i}\sim N(0,1)$ so that $\operatorname{Corr}(S_{i},\psi_{i})=r$ can be varied. We estimate $\operatorname{Cov}(\psi_{[n,i]},\psi_{[n,i+1]})$ based on 10,000 Monte Carlo trials by averaging it across all $i$. Figure 1 illustrates the non-negative correlation property shown in Proposition 3 by plotting the estimated covariance between $\psi_{[n,i]}$ and $\psi_{[n,i+1]})$ against the estimated correlation between $S_{i}$ and $\psi_{i}$ after averaging across all $i$. As the correlation between the treatment prioritization score $S_{i}$ and the individual treatment effect becomes stronger, the magnitude of positive covariance between the sorted individual treatment effects increases. Even when the score is negatively correlated with the individual treatment effect, the correlation of sorted individual treatment effects remains positive.
Proposition 3 implies that a properly normalized version of $p\widehat{\Psi}_{n}(p)$ converges to a stochastic process with non-negatively correlated increments. This key insight enables us to bound the process with an appropriately rescaled Wiener process, which has independent increments. Specifically, we use Slepian’s lemma, which states that mean-zero Gaussian stochastic processes are more tightly concentrated when the increments have a greater covariance (see Lemma S3 in Appendix S3). This leads to the following construction of a uniform confidence band.
Theorem 1 (Uniform Minimum-area Confidence Band for the GATES Estimator)
Under Assumptions 15 and for a given level of significance $\alpha\in(0,1)$, we have,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\widehat{\Psi}_{n}(p)-\frac{\beta_{0}^{\ast}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}-\beta_{1}^{\ast}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(p))}\right)\geq 1-\alpha.$
where the definition of $\mathbb{V}(\widehat{\Psi}_{n}(p))$ is given in Proposition 1 and
$\{\beta_{0}^{\ast}(\alpha),\beta_{1}^{\ast}(\alpha)\}=\operatorname*{argmin}_{\beta_{0},\beta_{1}\in\mathbb{R}_{+}^{2}}\int^{1}_{0}\beta_{0}+\beta_{1}\sqrt{t}\ \mathrm{d}t\ \text{ subject to }\ \mathbb{P}\left(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\forall t\in[0,1]\right)\geq 1-\alpha.$ (8)
with $W(t)$ denotes the standard Wiener process.
Proof is given in Appendix S3. Similar to Proposition 3, the inequality is tight when the treatment prioritization score is completely uninformative about the CATE.
To solve Equation (8), note that the objective function is linear in $\beta_{0}$ and $\beta_{1}$ and the probability expression, i.e., $\mathbb{P}\left(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\forall t\in[0,1]\right)$, is monotonically increasing in these coefficients. Therefore, a modified binary search algorithm on $(\beta_{0},\beta_{1})$ on a box $(\beta_{0},\beta_{1})\in[0,\overline{\beta}_{0}]\times[0,\frac{3}{4}\overline{\beta}_{0}]$ suffices, where $\overline{\beta}_{0}=\operatorname*{argmin}_{\beta_{0}\in\mathbb{R}}\mathbb{P}(W(t)\leq\beta_{0}))\geq 1-\alpha$, and $\overline{\beta}_{0}$ can be found through a standard line search. An example algorithm that can be used to solve Equation (8) is presented as Algorithm 1 where we approximate the probabilities through Monte Carlo simulation.
Input: Upper bound for $\beta_{0}$, denoted by $\overline{\beta}_{0}$, step size $\epsilon$, Type I error probability $\alpha$
      Output: Optimal values of $\{\beta_{0}(\alpha),\beta_{1}(\alpha)\}$
1:$\{V,\beta_{1l},\beta_{1h}\}\leftarrow\{\infty,0,\frac{3}{4}\overline{\beta}_{0}\}$
2:for  $\beta_{0c}$ from 0 to $\overline{\beta}_{0}$ with step $\epsilon$ do
3:     $\beta_{1l}\leftarrow 0$
4:     while $\beta_{1h}-\beta_{1l}>\epsilon$ do
5:         $\beta_{1m}\leftarrow\frac{\beta_{1l}+\beta_{1h}}{2}$
6:         if $\mathbb{P}\left(W(t)\leq\beta_{0c}+\beta_{1m}\sqrt{t},\;\forall t\in[0,1]\right)>1-\alpha$ then
7:              $\beta_{1h}\leftarrow\beta_{1m}$
8:         else
9:              $\beta_{1l}\leftarrow\beta_{1m}$
10:         end if
11:     end while
12:     if $\int^{1}_{0}\beta_{0c}+\beta_{1m}\sqrt{t}\ \mathrm{d}t<V$ then
13:         $V\leftarrow\int^{1}_{0}\beta_{0c}+\sqrt{t}\beta_{1m}\ \mathrm{d}t$
14:         $\{\beta_{0}(\alpha),\beta_{1}(\alpha)\}\leftarrow\{\beta_{0c},\beta_{1m}\}$
15:     end if
16:end for
17:return $\{\beta_{0}(\alpha),\beta_{1}(\alpha)\}$
Algorithm 1 Bounding the Weiner process by minimizing the area.
The uniform confidence band given in Theorem 1 minimizes the area of the resulting confidence region by approximating it via a specific functional form $\beta_{0}+\sqrt{t}\beta_{1}$. This choice is inspired by Kendall et al. (2007) who find it to perform well for Brownian motion. We also consider a slight modification of this minimum-area confidence bands. Policymakers are rarely interested in extremely small $p$, as the sample size is too small to provide a conclusive evidence. Thus, we may only wish to minimize the area of confidence region over $p\in[p_{l},1]$ for some $p_{l}\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\hat{\Psi}_{n}(p)-\frac{\beta_{0}^{\ast}(\alpha;p_{l})}{p}\sqrt{\mathbb{V}(\hat{\Psi}_{n}(1))}-\beta_{1}^{\ast}(\alpha;p_{l})\sqrt{\mathbb{V}(\hat{\Psi}_{n}(p))}\right)\geq 1-\alpha,$
where
$\{\beta_{0}^{\ast}(\alpha;p_{l}),\beta_{1}^{\ast}(\alpha;p_{l})\}=\operatorname*{argmin}_{\beta_{0},\beta_{1}\in\mathbb{R}_{+}^{2}}\left\{\int^{1}_{p_{l}}\beta_{0}+\beta_{1}\sqrt{t}\ \mathrm{d}t\;\;\text{subject to}\;\;\mathbb{P}(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\;\forall t\in[0,1])\geq 1-\alpha\right\}.$ (9)
Figure 2: Two functional forms (Equations (9) and (10)) used to create an one-sided 95% uniform confidence band for Brownian motion. 20 sample paths of Brownian motion are plotted in gray for reference.
The above functional form is not unique, and many other valid choices of confidence bands exist. As an example, we next extend Theorem 1 to provide confidence bands using a function of the form $t^{k}$ for any $20\leq k<1/$ 2. Figure 2 illustrates these two functional forms used for the confidence bands of Brownian motion.
Corollary 1
Under Assumptions 15, we have, for a given level of significance $\alpha$, and any $20\leq k<1/$ 2,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\widehat{\Psi}_{n}(p)-\gamma^{\ast}(\alpha;k)\mathbb{V}(\widehat{\Psi}_{n}(p))^{k}\mathbb{V}(\widehat{\Psi}_{n}(1))^{1/2-k}p^{2k-1}\right)\geq 1-\alpha.$
where
$\gamma^{\ast}(\alpha;k)=\inf\{\gamma\in\mathbb{R}\mid\mathbb{P}(W(t)\leq\gamma t^{k},\;\forall\;t\in[0,1])\geq 1-\alpha\}.$ (10)
Proof is given in Appendix S4.
Note that $k$ can only take values less than $1/2$. This is because the law of iterated logarithm states,
$\limsup_{t\to 0}\frac{W(t)}{\sqrt{2t\log\log\frac{1}{t}}}\to 1$
which implies that, for any $\gamma>0$, $\mathbb{P}(W(t)\leq\gamma\sqrt{t},\;\forall\;t\in[0,1])=0$. Thus, $\gamma^{\ast}(\alpha;k)$ does not exist for any $k\geq 1/2$. Since heuristically we expect $\mathbb{V}(\hat{\Psi}_{n}(p))$ to scale $O\left(p^{-1}\right)$, the width of the confidence bands for a given $k$ should scale roughly as $O\left(p^{k-1}\right)$. Therefore, if a policymaker is interested in selecting exceptional responders for a relatively small value of $p$, one should choose $k$ as close to $1/2$ as possible.
2.4 Numerical Illustration
We present a numerical example to illustrate the uniform confidence bands of different forms introduced above. We utilize the data generating process described in Section 3 and train the LASSO regression model by fitting it to the population data based on the empirical distribution of the full sample. Then, we generate a random sample of size $n=2,500$ from this population and apply our methodology to the CATE estimates from the LASSO regression. Specifically, we compute the proposed minimum-area uniform confidence bands in two ways, based on Equation (9) with $p_{l}=0$ and $p_{l}=0.05$, and Equation (10) with $k=0$ and $k=0.499$, using 10,000 Monte Carlo trials to approximate the probabilities. We also provide the pointwise confidence band as a reference.
(a) Entire range $p\in[0,1]$
(b) Around $p=0.15$
(c) Around $p=0.65$
Figure 3: Comparison of uniform minimum-area confidence bands using the sample of size $n=2500$.
Figure 2(a) shows the resulting confidence bands with the solid black line representing the point estimate $\widehat{\Psi}_{n}$, whereas Figures 2(b) and 2(c) zoom into the areas around $p=0.15$ and $p=0.65$, respectively. We find that in this numerical example, the confidence band based on Equation (10) with $k=0$ (solid orange line) has the greatest width when $p$ is small although it is slightly tighter than the other uniform minimum-area confidence bands when $p$ is large. The other three confidence bands are quite similar although the one based on Equation (10) with $k=0.499$ is slightly wider (approximately 5%) than those based on Equation (9).
We also observe that in general, the different types of uniform confidence bands are on average approximately $50\%$ wider than the pointwise confidence bands. This suggests that the uniform bands can be practically useful while retaining a strong statistical guarantee. The uniform bands, relative to the pointwise ones, are generally wider when $p$ becomes closer zero, and tighter when $p$ gets closer to one. This is because the variability of $\widehat{\Psi}(p)$ decreases as $p$ increases.
2.5 Extension to Other Evaluation Metrics
So far, we have focused on the identification of exceptional responders that maximizes $\Psi(p)=\mathbb{E}[\psi_{i}\mid F(S_{i})\geq p]$, using the individual-level evaluation metric, $\hat{\psi}_{i}=T_{i}Y_{i}/(n_{1}/n)-(1-T_{i})Y_{i}/(n_{0}/n)$. We emphasize, however, that Theorem 1 is applicable to any unit-level statistic,
$\hat{\psi}_{i}^{g}\ =\ g(Y_{i},T_{i},\mathbf{X}_{i},p),$
so long as the continuity and finite moment conditions (i.e., Assumptions 4 and 5) hold. This generalization allows for various alternative formulations of exceptional responder problem. For example, we may adjust for the cost of treatment and use the following alternative definition of $\Psi$,
$\Psi^{C}(p)\ =\ \mathbb{E}[Y_{i}(1)-Y_{i}(0)-c_{p}(\mathbf{X}_{i})\mid F(S_{i})\geq p]$
where $c_{p}(\cdot)$ is a known cost function based on the individual characteristics $\mathbf{X}_{i}$ and the proportion of exceptional responders $p$. Since this still leads to the individual level statistic $\hat{\psi}_{i}^{C}=\hat{\psi}_{i}-c_{p}(\mathbf{X}_{i})$, Theorem 1 is directly applicable.
Next, we consider the mean adjusted individual-level statistic, which can be written as,
$\hat{\psi}_{i}^{M}(\{Y_{i},T_{i},\mathbf{X}_{i}\}_{i=1}^{n},p)\ =\ h(Y_{i},T_{i},\mathbf{X}_{i},p)-\frac{p}{n}\sum_{i^{\prime}=1}^{n}h(Y_{i^{\prime}},T_{i^{\prime}},\mathbf{X}_{i^{\prime}},p),$
where $h(Y_{i},T_{i},\mathbf{X}_{i},p)$ is a unit-level statistic. Examples of such mean-adjusted statistics include the rank average treatment effect (RATE) (Yadlowsky et al., 2021) and the population average prescriptive effect (PAPE) (Imai and Li, 2023).
Here, motivated by the idea of the PAPE, we consider the following objective that adjusts for the overall average treatment,
$p^{\ast}\ =\ \operatornamewithlimits{argmax}_{p\in[0,1]}\Psi^{M}(p)\quad\text{where }\ \Psi^{M}(p)\ =\ \mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid F(S_{i})\geq p]-p\mathbb{E}[Y_{i}(1)-Y_{i}(0)].$
Then, the unit-level statistic is given by,
$\hat{\psi}_{i}^{M}(\{Y_{i},T_{i},\mathbf{X}_{i}\}_{i=1}^{n},p)\ =\ \frac{Y_{i}T_{i}}{n_{1}/n}-\frac{Y_{i}(1-T_{i})}{n_{0}/n}-\left(\frac{p}{n_{1}}\sum_{i=1}^{n}Y_{i}T_{i}-\frac{p}{n_{0}}\sum_{i=1}^{n}Y_{i}(1-T_{i})\right).$
The sample estimator of $\Psi^{M}(p)$ can be written as follows,
$\widehat{\Psi}^{M}_{n}(p)\ =\ \frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}^{M}_{[n,i]}\ =\ \frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]}-\frac{p\lfloor np\rfloor}{np}\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{[n,i]}=\widehat{\Psi}_{n}(p)-\frac{\lfloor np\rfloor}{n}\widehat{\Psi}_{n}(1).$ (11)
This alternative expression shows that $\widehat{\Psi}^{M}_{n}(p)$ can be constructed from $\widehat{\Psi}_{n}(p)$ in the same fashion as how a Brownian bridge on $[0,1]$ is constructed from a Wiener process. Therefore, Theorem 1 can be extended to the mean-adjusted statistic by replacing the Wiener process with the Brownian bridge.
Proposition 4
Under Assumptions 15, we have, for a given level of significance $\alpha\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\widehat{\Psi}^{M}_{n}(p)-\frac{\delta_{0}^{\ast}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}-\delta_{1}^{\ast}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(p))\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)}\right)\geq 1-\alpha,$
where
$\{\delta_{0}^{\ast}(\alpha),\delta_{1}^{\ast}(\alpha)\}=\operatorname*{argmin}_{\delta_{0},\delta_{1}\in\mathbb{R}}\left\{\int^{1}_{0}\delta_{0}+\delta_{1}\sqrt{t(1-t)}\ \mathrm{d}t\ \;\;\text{s.t.}\;\;\ \mathbb{P}\left(B(t)\leq\delta_{0}+\delta_{1}\sqrt{t(1-t)},\;\forall\;t\in[0,1]\right)\geq 1-\alpha\right\},$
with $B(t)$ denoting a standard Brownian Bridge on $[0,1]$.
Appendix S5 presents a proof of this proportion, which exploits the fact that Brownian Bridge $B(t)$ must be 0 at both $t=0$ and $t=1$.
3 A Simulation Study
We conduct a simulation study to examine the finite sample performance of the proposed methodology. We utilize one data generating process from the 2016 Atlantic Causal Inference Conference (ACIC) Competition. We briefly describe its simulation setting here and refer interested readers to Dorie et al. (2019) for additional details. The focus of this competition was the inference of the average treatment effect in observational studies. There are a total of 58 pre-treatment covariates, including 3 categorical, 5 binary, 27 count data, and 13 continuous variables. The data were taken from a real-world study with the sample size $n=4802$.
In our simulation, we assume that the empirical distribution of these covariates represent the population of interest and obtain each simulation sample via bootstrap. One important change from the original competition is that instead of utilizing a propensity model to determine $T$, we assume that the treatment assignment is completely randomized, i.e., $\Pr(T_{i}=1)=1/2$, and the treatment and control groups are of equal size, i.e., $n_{1}=n_{0}=n/2$.
To generate the outcome variable, we use one of the data generating processes used in the competition (the 28th), which is based on the generalized additive model with polynomial basis functions. The model represents a setting, in which there exists a substantial amount of treatment effect heterogeneity. For the sake of completeness, the formula for this outcome model is reproduced here:
$\mathbb{E}(Y_{i}(t)\mid\mathbf{X}_{i})$ $=1.60+0.53\times x_{29}-3.80\times x_{29}(x_{29}-0.98)(x_{29}+0.86)-0.32\times\mathbf{1}\{x_{17}>0\}$
$+0.21\times\mathbf{1}\{x_{42}>0\}-0.63\times x_{27}+4.68\times\mathbf{1}\{x_{27}<-0.61\}-0.39\times(x_{27}+0.91)\mathbf{1}\{x_{27}<-0.91\}$
$+0.75\times\mathbf{1}\{x_{30}\leq 0\}-1.22\times\mathbf{1}\{x_{54}\leq 0\}+0.11\times x_{37}\mathbf{1}\{x_{4}\leq 0\}-0.71\times\mathbf{1}\{x_{17}\leq 0,t=0\}$
$-1.82\times\mathbf{1}\{x_{42}\leq 0,t=1\}+0.28\times\mathbf{1}\{x_{30}\leq 0,t=0\}$
+$+\{0.58\times x_{29}-9.42\times x_{29}(x_{29}-0.67)(x_{29}+0.34)\}\times\mathbf{1}\{t=1\}$
$+(0.44\times x_{27}-4.87\times\mathbf{1}\{x_{27}<-0.80\})\times\mathbf{1}\{t=0\}-2.54\times\mathbf{1}\{t=0,x_{54}\leq 0\}.$
for $t=0,1$ where the error term is distributed according to the standard normal distribution.
We estimate the CATE and use it as our treatment prioritization score to identify exceptional responders. For the estimation of the CATE, we use Bayesian Additive Regression Trees (BART) (see Chipman et al., 2010; Hill, 2011; Hahn et al., 2020) and Causal Forest (Wager and Athey, 2018), and LASSO (Tibshirani, 1996). The models are trained by fitting them to the population data (i.e., the empirical distribution of the full sample). The number of trees were tuned through the 5-fold cross validation for both algorithms. For implementation, we use R 3.6.3 with the following packages: bartMachine (version 1.2.6) for BART, grf (version 2.0.1) for Causal Forest, and glmnet (version 4.1-2) for LASSO. The number of trees was tuned through 5-fold cross-validation for BART and Causal Forest. The regularization parameter was tuned similarly for LASSO.
ML algorithm Uniform Pointwise Pointwise×$\times 1.5$
$\bm{n=100}$ $\bm{n=500}$ $\bm{n=2500}$ $\bm{n=100}$ $\bm{n=500}$ $\bm{n=2500}$ $\bm{n=100}$ $\bm{n=500}$ $\bm{n=2500}$
BART $96.1\%$ $96.0\%$ $95.2\%$ $87.2\%$ $76.5\%$ $70.3\%$ $96.5\%$ $96.7\%$ $97.1\%$
Causal Forest $96.0\%$ $95.3\%$ $95.7\%$ $83.7\%$ $77.1\%$ $71.9\%$ $96.2\%$ $97.2\%$ $98.0\%$
LASSO $95.8\%$ $95.6\%$ $95.6\%$ $84.1\%$ $76.0\%$ $69.8\%$ $96.5\%$ $97.4\%$ $97.0\%$
Table 1: Empirical coverage of 95% confidence bands. The table presents the empirical coverage of uniform 95% confidence bands as presented in Theorem 1, along with its comparison to the coverage of the 95% pointwise confidence bands and $50\%$ inflated pointwise confidence bands. Coverage is computed as the percentage of experiments in which the confidence bands contains the entire GATES curve.
We investigate the finite-sample performance of the uniform confidence bands as defined in Theorem 1. Table 1 shows the proportion of simulation experiments, in which the confidence band contains the entire curve of $\Psi^{\ast}(p)$. We find that our uniform confidence bands reach the nominal coverage for every dataset size down to $n=100$ while the pointwise confidence bands suffer from a significant undercoverage, especially as the sample size $n$ grows. Although our estimators are theoretically conservative, we see little evidence of significant over-coverage across all algorithms and data sizes.
Following our observations in Section 2.4, we also show the coverage of the pointwise confidence bands after inflating it by $50\%$. We observe that this simple linear adjustment to pointwise confidence bands appear to achieve adequate coverage, suggesting that this may serve a heuristic but reasonable approximation to the $95\%$ uniform confidence bands.
Figure 4: Comparison of the lengths of the 95% uniform and pointwise confidence bands. The treatment prioritization score in this simulation is constructed using LASSO. The horizontal axis is the quantile threshold value of the score.
Finally, Figure 4 compares the lengths of these uniform and pointwise confidence intervals where the treatment prioritization score is estimated using LASSO. In general, the uniform confidence interval is about $25-50\%$ wider than the pointwise confidence intervals across different sample sizes.
4 Empirical Application
We apply the proposed methodology to the data from a clinical trial of late-stage prostate cancer.
4.1 Setup
Byar and Green (1980) present the data from a randomized placebo-controlled clinical trial to estimate the effect of diethylstilbestrol on overall survival of patients. Three different dosage levels (0.2 mg, 1.0mg, or 5.0mg) were used in the trial that involved 502 patients with stage 3 or 4 prostate cancer. For each patient, the trial recorded their baseline health characteristics, common laboratory measurements, prior disease history, and detailed information on the current prostate cancer. The primary outcome is the number of months of total survival at the end of follow-up, which may have occurred at either death or the program completion.
We will focus on the 5.0mg subgroup and consider the binary treatment of giving either 5.0mg estrogen or placebo for these patients. In total, 127 patients received the placebo while 125 patients received the estrogen treatment. The average treatment effect was estimated to be $-0.3$ months but this estimate was not statistically significant, suggesting that on average the estrogen treatment do not improve total survival.
By analyzing this data, we wish to identify a subgroup that may significantly benefit from the treatment although the overall average treatment effect estimate was not statistically significant. We utilize a total of 13 pre-treatment covariates, including age, weight, height, blood pressure, tumor size and stage, ECG diagnosis, and blood measurements. The estrogen treatment is denoted as $T_{i}=1$ whereas the placebo is represented as $T_{i}=0$. The outcome variable $Y_{i}$ is the total survival.
We randomly selected approximately 40% of the sample (i.e., 103 observations) as the training data and the reminder of the sample (i.e., 149 observations) as the evaluation data. We train the same three machine learning models as the ones used in our simulation study (Section 3). For Causal Forest, we tuned the number of trees on a grid of $[500,700,900,1100,1300,1500]$. For BART, tuning was done on the number of trees on a grid of $[50,70,90,110,130,150]$. For LASSO, we tuned the regularization parameter on a grid of $[0.01,0.05,0.1,0.5,1,5]$. All tuning was done with a 5-fold cross-validation procedure on the training set with mean-squared error as the validation metric.
4.2 Findings
Figure 5: Evaluation of different machine learning algorithms on the randomized clinical trial for Diethylstilbestrol. The dark blue solid line represents the point estimates of the number of months of the increase in total survival when maximally $p\%$ of the population is treated with Diethylstilbestrol. The dotted red line represents the pointwise $90\%$ confidence interval while the light blue dashed line represents the uniform $90\%$ confidence interval. The black horizontal line represents the average treatment effect, which is $-0.3$ months.
Figure 5 shows the results. For each plot, we compute the point estimates of the GATES (solid blue lines), the 90% minimum-area lower confidence bands (blue dotted lines) as shown in Equation (8), and 90% pointwise confidence bands (red dotted lines). We find that in this application, BART and especially Causal Forest appear to be able to identify some patterns of treatment effect heterogeneity in the population. In contrast, LASSO is unable to capture any statistically significant heterogeneity. We also find that the minimum-area uniform confidence bands are close to the pointwise bands for a larger value of $p$ and only diverge significantly for a small value of $p$.
Estimated proportion of Estimated 90% uniform
Estimator exceptional responders GATES confidence band
Causal Forest $18.8\%$ $27.2$ (2.27, $\infty$)
BART $32.2\%$ $18.1$ (0.53, $\infty$)
LASSO $91.2\%$ $1.35$ (--6.99, $\infty$)
Table 2: Estimated proportion of exceptional responders, their GATES estimates, and 90% uniform confidence bands. Each row represents the results based on different machine learning algorithms used to estimate the treatment prioritization score. The confidence intervals are corrected for multiple-testing using the Bonferroni method. The outcome is total survival in months.
Table 2 presents the estimated proportion of exceptional responders, their GATES estimates, and the corresponding 90% uniform confidence bands, for each ML algorithm used to construct the treatment prioritization score. The results show that both BART and Causal Forest appears to be able to identify a subset of population that would experience a statistically significant positive effect of the treatment on total survival. Specifically, we find that among these subsets of patients, 5mg of estrogen will increase their total survival by more than 15 months on average. In contrast, LASSO is unable to find a group of individuals who exhibit a statistically significantly positive effect.
Causal Forest in particular finds that $18\%$ of the population that would exhibit an increase of 27 months in total survival on average with the 90% uniform lower confidence band of 4.45 months. In comparison, the previous best optimization approach (e.g. Bertsimas et al., 2019) found a similarly sized group that could prolong total survival by 18 months on average but without a statistical guarantee. Although the exact subgroup might vary over multiple train-test splits, this example demonstrates that our methodology is able to exploit the power of modern machine learning algorithms while having a rigorous performance guarantee.
5 Concluding Remarks
As the use of machine learning tools become widespread in high-stake decision making, it is essential to deploy them effectively and safely. Doing so requires the accurate quantification of statistical uncertainty and the robust evaluation of their empirical performance. The proposed methodology provides a statistical performance guarantee when selecting an optimal subset of a target population who are predicted to benefit most from a treatment of interest. We believe that our methodology is particularly useful when the evaluation sample is of limited size, machine learning models provide biased or noisy estimates of treatment efficacy, and/or the proportion of exceptional responders is relatively small. In future research, we plan to extend the proposed methodology to other settings, including non-binary and dynamic treatments. It is also of interest to consider the question of how to account for the statistical uncertainty that arises from the estimation of treatment prioritization score, which is taken as given in our proposed framework.
Supplementary Appendix
Appendix S1 Proof of Proposition 2
We begin by proving two lemmas and stating Donsker’s Invariance Principle, which we use to prove Proposition 2.
Lemma S1
Under Assumptions 14 and for some $\epsilon>0$,
$\lim_{n\to\infty}\sup_{p\in[0,1]}n^{1/2+\epsilon}\left(\mathbb{E}[p\widehat{\Psi}_{n}(p)]-p\Psi(p)\right)=0.$
Proof.
We bound the bias of $\mathbb{E}(\widehat{\Psi}_{n}(p))$ by appealing to Proposition 1, which implies,
$|p\mathbb{E}(\widehat{\Psi}_{n}(p))-p\Psi(p)|\leq$ $\left|\mathbb{E}\left[\int_{F(c(p))}^{F(\hat{c}_{n}(p))}\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(x))\mathrm{d}x\right]\right|$ (S1)
By the definition of $\hat{c}_{n}(p)$, $F(\hat{c}_{n}(p))$ is the $np$th order statistic of $n$ independent uniform random variables. Therefore, it follows the Beta distribution with the shape and scale parameters equal to $np$ and $n-np+1$, respectively. Now, by Assumption 4, we can compute the first-order Taylor expansion of $\int^{x}_{a}\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(x))\mathrm{d}x$ and thus:
$\sup_{p\in[0,1]}|p\mathbb{E}(\widehat{\Psi}_{n}(p))-p\Psi(p)|\leq$ $\sup_{p\in[0,1]}\left|\mathbb{E}\left[\sup_{p\in[0,1]}\left|\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(p))\right|\{F(\hat{c}_{n}(p))-F(c(p))\}\right.\right.$
..$\biggl{.}\biggl{.}\hskip 72.26999pt+o(F(\hat{c}_{n}(p))-F(c(p)))\biggr{]}\biggr{|}$
$=$ $\sup_{p\in[0,1]}\left|\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(p))\right|\left|\frac{np}{n+1}-p\right|+o\left(\frac{1}{n}\right)$
$=$ $O\left(\frac{1}{n}\right).$
In particular, we thus have, for any $0<\epsilon<\frac{1}{2}$:
$\lim_{n\to\infty}n^{\frac{1}{2}+\epsilon}\sup_{p\in[0,1]}\left(\mathbb{E}[p\widehat{\Psi}_{n}(p)]-p\Psi(p)\right)=0.$
Lemma S2
Under Assumptions 14, we have for some $\epsilon>0$:
$\lim_{n\to\infty}n^{1/2+\epsilon}\sup_{p\in[0,1]}\int^{p}_{\frac{\lfloor np\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]\;\mathrm{d}x=0$
Proof.
By definition of the floor function, we have that $p-\frac{\lfloor np\rfloor}{n}\leq\frac{1}{n}$. Therefore for some $\epsilon<\frac{1}{2}$:
[…]
By Assumption 4, we know that $\sup_{p\in[0,1]}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=p]$ is finite. Therefore, the right handside of the above equation is zero, implying the desired result. ∎
Theorem S1 (Donsker’s Invariance Principle)
Consider a double sequence of random variables $U_{i,j}$ with $\mathbb{E}[U_{i,j}]=0$ for all $i,j\in\{1,2,\ldots,n\}$. Assume that $U_{n,i}$ satisfies the metric entropy integrability condition in $L_{2}$ as defined in Equation (3.1) of Ossiander (1987). Then, define the quantity:
$S_{n}(t)=\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}U_{n,i}$
where $t\in[0,1]$ and $S_{n}(0)=0$. Construct the random continuous polygon $s_{n}$ with basic points:
$b_{i}\ =\ \left(\frac{\mathbb{V}(S_{n}(i/n))}{\mathbb{V}(S_{n}(1))},\frac{S_{n}(i/n)}{\sqrt{\mathbb{V}(S_{n}(1))}}\right),$
where $i\in\{1,\cdots,n\}$. Then, as the sample size $n$ goes to infinity, $s_{n}$ converges in distribution to $G$ uniformly in the Skorokhod space, where $G$ is a zero-mean Gaussian process with covariance function
$K(p,p^{\prime})=\lim_{n\to\infty}\operatorname{Cov}\left(\frac{S_{n}(\lfloor pn\rfloor/n)}{\sqrt{\mathbb{V}(S_{n}(\lfloor pn\rfloor/n))}},\frac{S_{n}(\lfloor p^{\prime}n\rfloor/n)}{\sqrt{\mathbb{V}(S_{n}(\lfloor p^{\prime}n\rfloor/n))}}\right),$
in particular $K(p,p)=p$. Additionally, when $\operatorname{Cov}(U_{n,i},U_{n,j})=0$ for all $n$ and $1\leq i,j\leq n$, then $G$ is the standard Wiener process.
Finally, using the above results, we prove Proposition 2. Define for all $i\in\{1,\cdots,n\}$:
$\Delta\Psi_{n,i}=\frac{i}{n}\mathbb{E}\left[\hat{\psi}_{i}\ \biggl{|}\ F(S_{i})\geq\frac{i}{n}\right]-\frac{i-1}{n}\mathbb{E}\left[\hat{\psi}_{i}\ \biggl{|}\ F(S_{i})\geq\frac{i-1}{n}\right].$
Note that:
$p\widehat{\Psi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]},\qquad p\Psi(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\Delta\Psi_{n,i}+\int^{p}_{\frac{\lfloor np\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]\;\mathrm{d}x$
By Lemmas S1 and S2, we can construct a new random variable $U_{n,i}$ with $\mathbb{E}(U_{n,i})=0$ for some fixed quantity $\epsilon_{n,i}>0$ in the following manner,
$U_{n,i}$ $=\Delta\Psi_{n,i}-\hat{\psi}_{[n,i]}-\epsilon_{n,i},\quad\text{where}\ \lim_{n\to\infty}n^{1/2+\epsilon}\sup_{t\in[0,1]}\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\epsilon_{n,i}=0.$ (S2)
By Assumption 5 $\psi_{n,i}$ have finite second moments, so the induced order statistics $\psi_{[n,i]}$ satisfy the metric entropy integrability condition on $L_{2}$ Davydov and Egorov (2000). Therefore we can apply Theorem S1 to $U_{n,i}$ by considering a random continuous polygon $\tilde{s}_{n}$ connected by the basic points:
$\tilde{b}_{i}\ =\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\hat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})-\frac{1}{n}\sum_{i^{\prime}=1}^{i}\epsilon_{n,i^{\prime}}}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right),$
implying that $\tilde{s}_{n}$ converges to a zero-mean Gaussian Process. Equation (S2) implies $\frac{\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\epsilon_{n,i}}{\sqrt{\mathbb{V}(\Psi(1))}}\to 0$ uniformly in $t$. Thus, the random continuous polygon $s_{n}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right),$
converges to a zero-mean Gaussian process $G$. ∎
Appendix S2 Proof of Proposition 3
Proof.
Define $\psi(s)=\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=s)$. Then, we have,
$\operatorname{Cov}(\hat{\psi}_{[n,i]},\hat{\psi}_{[n,j]})$
=\ = $\operatorname{Cov}(\mathbb{E}[\hat{\psi}_{[n,i]}\mid S_{[n,i]}],\mathbb{E}[\hat{\psi}_{[n,j]}\mid S_{[n,j]}])+\mathbb{E}[\operatorname{Cov}(\hat{\psi}_{[n,i]},\hat{\psi}_{[n,j]}\mid S_{[n,i]},S_{[n,j]})]$
=\ = $\operatorname{Cov}(\psi(S_{[n,i]}),\psi(S_{[n,j]}))$
where $S_{[n,i]}$ is the $i$th order statistic of $S_{i}$ given the sample of size $n$. We follow the proof strategy of (Cuadras, 2002, Theorem 1) by defining $(S_{[n,i]}^{\prime},S_{[n,j]}^{\prime})$ as random variables that are independently and identically distributed as $(S_{[n,i]},S_{[n,j]})$. Then we have:
$\operatorname{Cov}(\psi(S_{[n,i]}),\psi(S_{[n,j]}))$ $=\mathbb{E}[\psi(S_{[n,i]})\psi(S_{[n,j]})]-\mathbb{E}[\psi(S_{[n,i]})]\mathbb{E}[\psi(S_{[n,j]})]$
$=\frac{1}{2}\mathbb{E}[(\psi(S_{[n,i]})-\psi(S_{[n,i]}^{\prime}))(\psi(S_{[n,j]})-\psi(S_{[n,j]}^{\prime}))]$
Since $\int_{-\infty}^{\infty}\mathbbm{1}(a\leq S_{[n,i]})-\mathbbm{1}(a\leq S_{[n,i]}^{\prime})\mathrm{d}\psi(a)=\psi(S_{[n,i]})-\psi(S_{[n,i]}^{\prime})$, we have:
$\operatorname{Cov}(\psi(S_{[n,i]}),\psi(S_{[n,j]}))$
$=$ $\frac{1}{2}\mathbb{E}\left[\int_{R^{2}}\left\{\mathbbm{1}(a\leq S_{[n,i]})-\mathbbm{1}(a\leq S_{[n,i]}^{\prime})\right\}\left\{\mathbbm{1}(b\leq S_{[n,j]})-\mathbbm{1}(b\leq S_{[n,j]}^{\prime})\right\}\mathrm{d}\psi(a)\mathrm{d}\psi(b)\right].$
Under the assumptions, $\psi(\cdot)$ is a continuous function and $\psi(S_{[n,i]})$ has a finite second moment. Therefore, we can apply Fubini’s theorem to exchange integration signs:
$\operatorname{Cov}(\psi(S_{[n,i]}),\psi(S_{[n,j]}))=$ $\int_{R^{2}}1-\mathbb{P}(S_{[n,i]}<a)\mathbb{P}(S_{[n,j]}<b)+\mathbb{P}(S_{[n,i]}<a,S_{[n,j]}<b)$
$\quad-(1-\mathbb{P}(S_{[n,i]}<a))(1-\mathbb{P}(S_{[n,j]}<b))\mathrm{d}\psi(a)\mathrm{d}\psi(b)$
$=$ $\int_{R^{2}}\mathbb{P}(S_{[n,i]}<a,S_{[n,j]}<b)-\mathbb{P}(S_{[n,i]}<a)\mathbb{P}(S_{[n,j]}<b)\mathrm{d}\psi(a)\mathrm{d}\psi(b).$
Since $S_{[n,i]}$ and $S_{[n,j]}$ are order statistics, by Theorem 3.4 of Boland et al. (1996), we know that $\mathbb{P}(S_{[n,i]}<a,S_{[n,j]}<b)\geq\mathbb{P}(S_{[n,i]}<a)\mathbb{P}(S_{[n,j]}<b)$ for all $a,b$. Thus, the desired conclusion holds. ∎
Appendix S3 Proof of Theorem 1
To prove Theorem 1, we first state Slepian’s lemma.
Lemma S3 (Slepian (1962))
For Gaussian random variables $X=(X_{1},\cdots,X_{n})$ and $Y=(Y_{1},\cdots,Y_{n})$ in $\mathbb{R}^{n}$ satisfying $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]=0$, $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]$ and $\mathbb{E}[X_{i}X_{j}]\leq\mathbb{E}[Y_{i}Y_{j}]$ for all $i,j=1,\cdots,n$, we have for any real numbers $u_{1},\cdots u_{n}$:
$\mathbb{P}\left[\bigcap_{i=1}^{n}\{X_{i}\geq u_{i}\}\right]\leq\mathbb{P}\left[\bigcap_{i=1}^{n}\{Y_{i}\geq u_{i}\}\right].$
We define a set of independent Gaussian variables $\phi_{[n,i]}$ whose mean and variance are identical to those of $\hat{\psi}_{[n,i]}$, i.e., $\mathbb{E}(\phi_{[n,i]})=\mathbb{E}(\hat{\psi}_{[n,i]})$, $\mathbb{V}(\phi_{[n,i]})=\mathbb{V}(\hat{\psi}_{[n,i]})$, and $\operatorname{Cov}(\phi_{[n,i]},\phi_{[n,j]})=0$. Then, define:
$p\widehat{\Phi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\phi_{[n,i]}.$
Since $\phi_{[n,i]}$ are independent, the metric entropy integrability condition on $L_{2}$ is automatically satisfied by Assumption 5. Applying Theorem S1, the random continuous polygon $s_{n}^{\prime}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Phi}_{n}(1))},\frac{\frac{i}{n}\Phi(\frac{i}{n})-\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Phi}_{n}(1))}}\right)$
converges to $W$, where $W$ is the standard Wiener process. Then, we have:
[…]
where
$\{\beta_{0}^{\ast}(\alpha),\beta_{1}^{\ast}(\alpha)\}=\operatorname*{argmin}_{\beta_{0},\beta_{1}\in\mathbb{R}_{+}^{2}}\left\{\int^{1}_{0}\beta_{0}+\beta_{1}\sqrt{t}\ \mathrm{d}t\;\;\text{subject to}\;\;\mathbb{P}(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\forall\;t\in[0,1])\geq 1-\alpha\right\}.$
Proposition 3 implies that the covariance function $K_{G}$ of the Gaussian Process $G$ satisfies $K_{G}(p,p^{\prime})\geq K_{W}(p,p^{\prime})=0$ where $K_{W}$ is the covariance function of $W$. In addition, we have $K_{G}(p,p)=K_{W}(p,p)$ by construction. Thus, we can apply Slepian’s lemma (Lemma S3) to Equation (S3) and obtain,
$\mathbb{P}\left(G(p)\geq-\beta^{*}_{0}(\alpha)-\beta^{*}_{1}(\alpha)\sqrt{p},\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$ (S4)
Proposition 2 implies that uniformly for all $p\in[0,1]$, we have as $n\to\infty$:
$\left(\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right)\to(p,G(p))$
Therefore, we can rewrite Equation (S4) as:
$\lim_{n\to\infty}\mathbb{P}\left(\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\geq-\beta^{*}_{0}(\alpha)-\sqrt{\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\beta^{*}_{1}(\alpha),\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$
Rearranging the terms gives:
$\lim_{n\to\infty}\mathbb{P}\left(p\Psi(p)\geq p\widehat{\Psi}_{n}(p)-\beta^{*}_{0}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}-\beta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(p\widehat{\Psi}_{n}(p))},\;\forall p\in[0,1]\right)\geq 1-\alpha.$
Therefore, for any $c_{0}>0$, we have:
$\lim_{n\to\infty}\mathbb{P}\left(\Psi(p)\geq\widehat{\Psi}_{n}(p)-\frac{\beta^{*}_{0}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}-\beta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(p))},\;\forall\;p\in[c_{0},1]\right)\geq 1-\alpha.$
Taking $c_{0}\to 0$ gives the required result. ∎
Appendix S4 Proof of Corollary 1
Identical to Theorem 1, we define a set of independent Gaussian variables $\phi_{[n,i]}$ whose mean and variance are identical to those of $\hat{\psi}_{[n,i]}$, i.e., $\mathbb{E}(\phi_{[n,i]})=\mathbb{E}(\hat{\psi}_{[n,i]})$, $\mathbb{V}(\phi_{[n,i]})=\mathbb{V}(\hat{\psi}_{[n,i]})$, and $\operatorname{Cov}(\phi_{[n,i]},\phi_{[n,j]})=0$. Then, define:
$p\widehat{\Phi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\phi_{[n,i]}.$
Since $\phi_{[n,i]}$ are independent, the metric entropy integrability condition on $L_{2}$ is automatically satisfied by Assumption 5. Applying Theorem S1, the random continuous polygon $s_{n}^{\prime}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Phi}_{n}(1))},\frac{\frac{i}{n}\Phi(\frac{i}{n})-\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Phi}_{n}(1))}}\right)$
converges to $W$, where $W$ is the standard Wiener process. Then, we have:
[…]
where
$\gamma^{\ast}(\alpha;k)=\inf\{\gamma\in\mathbb{R}\mid\mathbb{P}(W(t)\leq\gamma t^{k},\;\forall\;t\in[0,1])\geq 1-\alpha\}.$
Proposition 3 implies that the covariance function $K_{G}$ of the Gaussian Process $G$ satisfies $K_{G}(t,t^{\prime})\geq K_{W}(t,t^{\prime})=0$ where $K_{W}$ is the covariance function of $W$. In addition, we have $K_{G}(t,t)=K_{W}(t,t)$ by construction. Thus, we can apply Slepian’s lemma (Lemma S3) to Equation (S3) and obtain,
$\mathbb{P}\left(G(t)\geq-\gamma^{\ast}(\alpha;k)t^{k},\;\forall\;t\in[0,1]\right)\geq 1-\alpha.$ (S6)
Proposition 2 implies that uniformly for all $p\in[0,1]$, we have as $n\to\infty$:
$\left(\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right)\to(p,G(p))$
Therefore, we can rewrite Equation (S6) as:
$\lim_{n\to\infty}\mathbb{P}\left(\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\geq-\gamma^{\ast}(\alpha;k)\left(\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))}\right)^{k},\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$
Rearranging the terms gives:
$\lim_{n\to\infty}\mathbb{P}\left(p\Psi(p)\geq p\widehat{\Psi}_{n}(p)-\gamma^{\ast}(\alpha;k)p^{2k}\mathbb{V}(\widehat{\Psi}_{n}(p))^{k}\mathbb{V}(\widehat{\Psi}_{n}(1))^{1/2-k},\;\forall p\in[0,1]\right)\geq 1-\alpha.$
Therefore, for any $c_{0}>0$, we have:
$\lim_{n\to\infty}\mathbb{P}\left(\Psi(p)\geq\widehat{\Psi}_{n}(p)-\gamma(\alpha;k)^{\ast}p^{2k-1}\mathbb{V}(\widehat{\Psi}_{n}(p))^{k}\mathbb{V}(\widehat{\Psi}_{n}(1))^{1/2-k},\;\forall\;p\in[c_{0},1]\right)\geq 1-\alpha.$
Taking $c_{0}\to 0$ gives the required result. ∎
Appendix S5 Proof of Proposition 4
First, we introduce Kahane’s Inequality.
Lemma S4 (Kahane (1985)’s Inequality)
Consider Gaussian random variables $X=(X_{1},\ldots,X_{n})$ and $Y=(Y_{1},\ldots,Y_{n})$ in $\mathbb{R}^{n}$ such that $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]=0$, $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]$, and $\operatorname{Cov}(X_{i},X_{j})\leq\operatorname{Cov}(Y_{i},Y_{j})$ for all $i,j=1,\cdots,n$. Define $Z=(Z_{1},\ldots,Z_{n})$ where $Z_{i}=\sqrt{1-t^{2}}X_{i}+tY_{i}$ for $t\in[0,1]$. If a twice-differentiable function function $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfies for all $i\neq j$ and for all $t\in[0,1]$, i.e., $\mathbb{E}\left[\frac{\partial^{2}f(Z)}{\partial Z_{i}\partial Z_{j}}\right]\geq 0$. Then, we have $\mathbb{E}[f(X)]\leq\mathbb{E}[f(Y)]$.
We then utilize Lemma S4 to prove the following useful “conditional” Slepian inequality.
Lemma S5
For Gaussian random variables $X=(X_{1},\ldots,X_{n})$ and $Y=(Y_{1},\ldots,Y_{n})$ in $\mathbb{R}^{n}$ satisfying $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]=0$, $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]$ and $0\leq\operatorname{Cov}(X_{i},X_{j})\leq\operatorname{Cov}(Y_{i},Y_{j})$ for all $i,j=1,\cdots,n$, we have for any $u_{1},\ldots u_{n}\leq 0$:
$\mathbb{P}\left[\bigcap_{i=1}^{n-1}\{X_{i}\geq u_{i}\}\ \Bigl{|}\ X_{n}=0\right]\leq\mathbb{P}\left[\bigcap_{i=1}^{n-1}\{Y_{i}\geq u_{i}\}\ \Bigl{|}\ Y_{n}=0\right].$
Proof.
Consider the function
$f(x_{1},\ldots,x_{n})=\prod_{i=1}^{n-1}\mathbf{1}(x_{i}\geq u_{i})\times\mathbf{1}(x_{n}=0)$
Our goal is to prove $\mathbb{E}[f(X)]\leq\mathbb{E}[f(Y)]$, which immediately implies the desired result. Unfortunately, $f$ is not twice-differentiable, so we approximate it by the following infinitely differentiable function as $\epsilon>0$ approaches 0:
$f_{\epsilon}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n-1}\frac{1}{1+e^{-(x_{i}-u_{i}-\sqrt{\epsilon})/\epsilon}}\times e^{-x_{n}^{2}/\epsilon}$
By construction, we have $\frac{\partial^{2}f_{\epsilon}(x_{1},\ldots,x_{n})}{\partial x_{i}\partial x_{j}}\geq 0$ for all $i,j\neq n$, and so we only need to consider $\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\right]$ where $Z=(Z_{1},\ldots,Z_{n})$ and $Z_{i}=\sqrt{1-t^{2}}X_{i}+tY_{i}$ for $t\in[0,1]$. By definition, we have $\operatorname{Cov}(Z_{i},Z_{j})\geq 0$. Since $u_{i}\leq 0$, we have:
$\mathbb{P}(Z_{i}\leq u_{i},Z_{n}<0)>\mathbb{P}(Z_{i}\leq u_{i},Z_{n}>0)$ (S7)
Now note that the function $f_{\epsilon}(x_{1},\ldots,x_{n})$ is symmetric around zero for $x_{n}$ and approaches the indicator function $\mathbf{1}(x_{i}\geq u_{i})$ from the left for $x_{i}$ satisfying $i\neq n$. Therefore, as $\epsilon$ approaches 0 and $i\neq n$, we have,
$\lim_{\epsilon\downarrow 0}\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\right]$ $=$ $\mathbb{P}(Z_{i}\leq u_{i})\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\ \biggl{|}\ Z_{i}\leq u_{i}\right]$
$=$ $\ \mathbb{P}(Z_{i}\leq u_{i},Z_{n}>0)\underbrace{\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\ \biggl{|}\ Z_{i}\leq u_{i},Z_{n}>0\right]}_{<0}$
$+\mathbb{P}(Z_{i}\leq u_{i},Z_{n}<0)\underbrace{\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\ \biggl{|}\ Z_{i}\leq u_{i},Z_{n}<0\right]}_{>0}$
$>$ $0$
where the last inequality follows from Equation (S7). Thus, the desired result follows immediately from the application of Lemma S4 to $f_{\epsilon}$ and letting $\epsilon$ apprach 00. ∎
Proposition 2 implies that the random continuous polygon connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right)$
converges to a zero-mean Gaussian process $G(t)$ such that $K_{G}(t,t^{\prime})\geq K_{W}(t,t^{\prime})=0$ where $K_{W}$ is the covariance function of a standard Weiner process $W$. Furthermore, Equation (11) and the continuous mapping theorem imply that the random continuous polygon connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))},\frac{\frac{i}{n}\Psi^{M}(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}}\right)$
converges to a zero-mean Gaussian bridge $G(t)-tG(1)$ on $[0,1]$. By the theory of Gaussian bridges (see e.g., Theorem 3.1 in Sottinen and Yazigi (2014)), we can equivalently characterize this Gaussian bridge as $G(t)\mid G(1)=0$.
Define a set of independent Gaussian variables $\phi_{[n,i]}$ whose mean and variance are identical to those of $\hat{\psi}_{[n,i]}$, i.e., $\mathbb{E}(\phi_{[n,i]})=\mathbb{E}(\hat{\psi}_{[n,i]})$, $\mathbb{V}(\phi_{[n,i]})=\mathbb{V}(\hat{\psi}_{[n,i]})$, and $\operatorname{Cov}(\phi_{[n,i]},\phi_{[n,j]})=0$. Then, define:
$\widehat{\Phi}_{n}(p)$ $=\frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\phi_{[n,i]},$
$\widehat{\Phi}^{M}_{n}(p)$ $=\ \widehat{\Phi}_{n}(p)-\frac{p\lfloor np\rfloor}{np}\widehat{\Phi}_{n}(1).$
Theorem S1 implies that the random continuous polygon connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Phi}^{M}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Phi}^{M}_{n}(1))},\frac{\frac{i}{n}\Phi^{M}(\frac{i}{n})-\frac{i}{n}\widehat{\Phi}^{M}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Phi}^{M}_{n}(1))}}\right)$
converges to the Brownian Bridge $B(t)=W(t)-tW(1)$. Similarly, this Gaussian process can be characterized as $W(t)\mid W(1)=0$. Then, we have:
[…]
where
$\{\delta^{*}_{0}(\alpha),\delta^{*}_{1}(\alpha)\}$
$=\operatorname*{argmin}_{\delta_{0},\delta_{1}\in\mathbb{R}_{+}^{2}}\left\{\int^{1}_{0}\delta_{0}+\delta_{1}\sqrt{t(1-t)}\ \mathrm{d}t\;\;\text{subject to}\;\;\mathbb{P}(W(t)\leq\delta_{0}+\delta_{1}\sqrt{t(1-t)},\;\forall\;t\in[0,1]\mid W(1)=0)\geq 1-\alpha\right\}.$
The application of Lemma S5 to the stochastic processes $G(t)$ and $W(t)$ implies that we can rewrite Equation (S8) as:
$\mathbb{P}\left(G(t)\geq-\delta^{*}_{0}(\alpha)-\delta^{*}_{1}(\alpha)\sqrt{t(1-t)},\;\forall\;t\in[0,1]\mid G(1)=0\right)\geq 1-\alpha.$ (S9)
Proposition 2 implies that uniformly for all $t\in[0,1]$, we have as $n\to\infty$:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))},\frac{\frac{i}{n}\Psi^{M}(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}}\right)\to(t,G(t)\mid G(1)=0)$
Therefore, we can rewrite Equation (S9) as:
$\lim_{n\to\infty}\mathbb{P}\left(\frac{p\Psi^{M}(p)-p\widehat{\Psi}^{M}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}}\geq-\delta^{*}_{0}(\alpha)-\delta^{*}_{1}(\alpha)\sqrt{\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)},\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$
Rearranging the terms gives:
$\lim_{n\to\infty}\mathbb{P}\left(p\Psi^{M}(p)\geq p\widehat{\Psi}^{M}_{n}(p)-\delta^{*}_{0}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}-\delta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)},\;\forall p\in[0,1]\right)\geq 1-\alpha.$
Therefore, for any $c_{0}>0$, we have:
$\lim_{n\to\infty}\mathbb{P}\left(\Psi^{M}(p)\geq\widehat{\Psi}^{M}_{n}(p)-\frac{\delta^{*}_{0}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}-\delta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(p))\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)},\;\forall\;p\in[c_{0},1]\right)\geq 1-\alpha.$
Taking $c_{0}\to 0$ gives the required result. ∎