### v3

Abstract
Across a wide array of disciplines, many researchers use machine learning (ML) algorithms to identify a subgroup of individuals who are likely to benefit from a treatment the most (“exceptional responders”) or those who are harmed by it. A common approach to this subgroup identification problem consists of two steps. First, researchers estimate the conditional average treatment effect (CATE) using an ML algorithm. Next, they use the estimated CATE to select those individuals who are predicted to be most affected by the treatment, either positively or negatively. Unfortunately, CATE estimates are often biased and noisy. In addition, utilizing the same data to both identify a subgroup and estimate its group average treatment effect results in a multiple testing problem. To address these challenges, we develop uniform confidence bands for estimation of the group average treatment effect sorted by generic ML algorithm (GATES). Using these uniform confidence bands, researchers can identify, with a statistical guarantee, a subgroup whose GATES exceeds a certain effect size, regardless of how this effect size is chosen. The validity of the proposed methodology depends solely on randomization of treatment and random sampling of units. Importantly, our method does not require modeling assumptions and avoids a computationally intensive resampling procedure. A simulation study shows that the proposed uniform confidence bands are reasonably informative and have an appropriate empirical coverage even when the sample size is as small as 100. We analyze a clinical trial of late-stage prostate cancer and find a relatively large proportion of exceptional responders.
Key Words: causal inference, exceptional responders, heterogeneous treatment effects, treatment prioritization, uniform confidence bands
1 Introduction
In a diverse set of fields, machine learning (ML) algorithms are frequently used to identify a subgroup of individuals who are likely to benefit from a treatment the most (so called “exceptional responders”) or those who are negatively affected by it. In clinical studies, for example, identifying those who experience surprisingly long-lasting positive health outcomes when a majority of patients do not, is useful for understanding why a treatment works and determining which patients should receive it (e.g., Takebe et al., 2015; Conley et al., 2020). Similarly, it is critical to identify a subpopulation of patients who are harmed by a commonly used treatment. In public policy, finding those individuals who are likely to be helped by a treatment is a key methodological step for improving targeting strategies (e.g., Imai and Ratkovic, 2013; Chernozhukov et al., 2019).
A widely adopted strategy for subgroup identification proceeds in two steps. First, an ML algorithm is used to estimate the conditional average treatment effect (CATE) or its proxy such as biomarker response, conditional on pre-treatment covariates. Second, individuals are ranked by the estimated score, and those in the top (or bottom) of the distribution are selected as exceptional responders (or those who are most negatively affected). For example, one might define the exceptional responders as those in the top quantile of the score and report the average treatment effect within this group. Such a two-step procedure inevitably compresses rich heterogeneity into a one-dimensional score and may therefore be suboptimal in identifying exceptional responders. Nevertheless, it is commonly used in clinical and policy settings due to its interpretability and practical utility (e.g., Sussman et al., 2013; Razonable et al., 2022; Imai et al., 2023). In the literature, the average treatment effect among individuals grouped by the prioritization score is referred to as the Grouped Average Treatment Effect or GATES (Chernozhukov et al., 2019).
As a motivating study, we consider a well-known randomized clinical trial described by Byar and Green (1980) concerning the effect of administering a type of synthetic estrogen, called diethylstilbestrol, to prolong total survival among late-stage prostate cancer patients. The trial was overall unsuccessful as the treatment group with the highest dosage, 5.0mg, registered an average of 0.3 month decrease in total survival compared to the placebo group. However, given that the biological pathway of estrogen in reducing prostate cancer growth is well-known, clinical researchers believed that at least a certain subgroup of patients could have benefited from the treatment. Therefore, many researchers have attempted to find such a subgroup that exhibits a large GATES, by first applying ML, optimization, or heuristic methods to create a treatment prioritization score and then optimize over the threshold of the score with a certain objective function (e.g. Bonetti and Gelber, 2000; Rosenkranz, 2016; Ballarini et al., 2018; Bertsimas et al., 2019).
Despite its intuitive appeal and frequent use in practice, this approach faces several methodological challenges. First, there exists a considerable degree of statistical uncertainty in estimating the GATES for the selected subgroup. The treatment prioritization scores are often based on noisy and imperfect estimates of the CATE or other target quantities. Second, the subgroup of exceptional responders (or those who are negatively affected by the treatment) is usually small, and to make the matter worse, the size of experimental data used to identify such individuals is often not large either. Third, accounting for this statistical uncertainty is difficult because utilizing the same data to both evaluate the GATES and select the group of potential exceptional responders suffers from a multiple testing problem. Although this problem can be alleviated by setting aside a large holdout set for evaluation, such an approach is often not a feasible solution given the limited sample size of randomized control trials.
In this paper, we develop uniform confidence bands (intervals) for estimating the GATES among a subgroup of individuals defined by a quantile of treatment prioritization score. This group represents the subgroup of individuals who are predicted to benefit most from (or be harmed most by) the treatment. The proposed uniform confidence bands are valid regardless of the way in which the quantile value is chosen, thereby providing a statistical guarantee for a generic objective function to be optimized when selecting the group of individuals who are either most positively or most negatively affected by the treatment. For example, our methodology enables researchers to choose the quantile cutoff such that the GATES among the selected exceptional responders exceeds a certain threshold with a pre-specified probability.
The proposed methodology is based on Neyman’s repeated sampling framework (Neyman, 1923), assuming only randomization of treatment and random sampling of individual units without requiring any modeling assumption or repeated sampling methods. Thus, our methodology is applicable to any generic treatment prioritization score, including the ones that may not have any performance guarantee. Although we do not pursue in this paper, additional assumptions, such as monotonicity of treatment effect in treatment prioritization score, can be imposed to further obtain tighter confidence bands at the expense of generality (e.g., Chernozhukov et al., 2010). Our simulation study shows that the proposed uniform confidence bands are reasonably informative and provide an adequate empirical coverage for sample sizes that are as small as $n=100$ (Section 3).
To develop our methodology (Section 2), we first show that the problem of finding uniform confidence bands is equivalent to deriving the distribution of a particular functional of induced order statistics. This functional is also called a concomitant of order statistics in the literature, and has been the subject of independent interest (e.g., David and Galambos, 1974; Yang, 1977; Sen, 1976). In particular, Bhattacharya (1974) and Davydov and Egorov (2000) explore the asymptotic properties of general functionals of induced order statistics.
While the resulting asymptotic distributions are in general difficult to characterize, we show that the GATES are non-negatively correlated with one another regardless of how the treatment prioritization score is constructed. This allows us to uniformly bound the GATES by a Wiener process through an application of Donsker’s theorem. Although the resulting uniform confidence bands are not unique, we propose a specific choice that yields a narrow band for a reasonable range of quantile values. We further extend our theoretical results beyond the GATES to other causal quantities of interest, widening the applicability of the proposed methodology.
Finally, we apply the proposed methodology to the aforementioned randomized clinical trial (Section 4). We find that, with 95% of confidence, modern ML algorithms are able to identify a subset of patients whose survival would be increased by the treatment up to 27 months on average. This contrasts with a previous study which found, albeit without a statistical guarantee, a group of exceptional responders whose survival is increased by only up to 18 months on average (e.g., Bertsimas et al., 2019). This empirical application demonstrates that our methodology can exploit the power of modern ML algorithms while providing a statistical performance guarantee.
Related Literature.
The proposed methodology is related to several strands of the existing literature in computer science, optimization, and statistics. First, there exists a large literature whose goal is the development of methods to identify subgroups with large treatment effects using ML algorithms (e.g. Kehl and Ulm, 2006; Su et al., 2009; Foster et al., 2011; Hardin et al., 2013; Luedtke and van der Laan, 2016; Bertsimas et al., 2019; Cho and Ghosh, 2021; Bonvini et al., 2023). However, these existing studies either provides no statistical performance guarantee or assumes that the threshold used for subgroup selection is pre-specified. In contrast, we focus on developing uniform confidence bands that are valid regardless of how such a threshold is chosen.
Another closely related literature is the one about maximally selected test statistics (e.g., Miller and Siegmund, 1982; Hothorn and Lausen, 2002). Cast into the setting of our paper, these authors develop a statistical test of the null hypothesis that the CATE is independent of the treatment prioritization score. While this mathematical formulation is similar, constructing uniform confidence bands must allow for arbitrary dependence. As discussed in Section 2.3, this requires us to use a novel proof strategy that is different from the ones used in the literature.
Third, Chernozhukov et al. (2019) considers the estimation of the GATES based on the data from randomized experiments. Although they use resampling methods to provide statistical inference, the resulting confidence bands are point-wise and are not uniformly valid across the quantile values. While Imai and Li (2022) improves upon this methodology by using Neyman’s repeated sampling framework (thereby, avoiding computationally intensive resampling methods), the authors also only provide point-wise confidence bands. Our work fills this methodological gap by developing uniform inference for the GATES estimation.
Finally, there exist a small but fast growing body of literature that develops statistical methods for experimentally evaluating the performance of data-driven individualized treatment rules (ITRs). Imai and Li (2023) use Neyman’s repeated sampling framework to experimentally evaluate the efficacy of ITRs. Building on the up-lift modeling literature (e.g., Radcliffe, 2007), the authors propose the Population Average Prescriptive Effect (PAPE) curve as a performance measure of an ITR relative to the non-individualized treatment rule that treats the same proportion of randomly selected individuals. This is a causal analogue of the Receiver Operating Characteristic curve. Like this paper, Imai and Li (2023) provide an assumption-free statistical inference methodology to evaluate the performance of ITRs. Unlike our proposed methodology, however, their confidence bands are point-wise and do not yield uniform inference for the entire curve.
Similarly, Yadlowsky et al. (2021) considers the performance evaluation of ITRs, using a related quantity called the Targeted Operating Characteristic (TOC) curve. The TOC curve is similar to the prescriptive effect curve but compares the performance of an ITR with the treatment rule that treats everyone. While the authors show how to quantify statistical uncertainty for the area under the TOC curve, their inferential methods are point-wise and are not uniformly valid for the entire TOC curve. Unlike these two existing works, the proposed methodology yields uniform statistical inference, enabling the selection of exceptional responders and other subgroups with uniform statistical guarantees. As shown below, the proposed uniform confidence bands can be applied to the PAPE and TOC curves as well.
2 The Proposed Methodology
2.1 Setup and Assumptions
Suppose that we have a total of $n$ units independently sampled from a (super)population $\mathcal{Q}$. We consider a classical randomized experiment where a binary treatment $T_{i}\in\{0,1\}$ is completely randomized with a total of $n_{1}$ units assigned to the treatment condition and the remaining $n_{0}=n-n_{1}$ units assigned to the control condition. For each unit $i=1,\ldots,n$, we observe the outcome $Y_{i}\in\mathcal{Y}$ as well as a vector of pre-treatment covariates, $\bm{X}_{i}\in\mathcal{X}$. The observed outcome for unit $i$ is given by $Y_{i}=T_{i}Y_{i}(1)+(1-T_{i})Y_{i}(0)$ where $Y_{i}(t)$ represents the potential outcome under the treatment condition $T_{i}=t$. This implies the stable unit treatment value assumption or SUTVA (Rubin, 1990). In particular, the outcome of one unit is assumed to be not affected by the treatment condition of other units. We summarize our setup here:
Assumption 1 (Random Sampling of Units)
We assume independently and identically distributed observations,
$\{\bm{X}_{i},Y_{i}(1),Y_{i}(0)\}_{i=1}^{n}\ \stackrel{{\scriptstyle\rm i.i.d.}}{{\sim}}\ \mathcal{Q}$
Assumption 2 (Random Assignment of Treatment)
We assume complete randomization of treatment assignment,
$\Pr(T_{i}=1\mid\{\bm{X}_{i},Y_{i^{\prime}}(1),Y_{i^{\prime}}(0)\}_{i=1}^{n})=\frac{n_{1}}{n}\quad\text{for all }i=1,\ldots,n,\ \text{and }n_{1}=\sum_{i=1}^{n}T_{i}$
Without the loss of generality, we assume that a positive treatment effect is desirable. We consider a scoring rule $f(\cdot)$ that determines treatment prioritization where a greater score represents a higher priority. Formally, a scoring rule maps the pre-treatment covariates to a scalar variable,
$f:\mathcal{X}\longrightarrow\mathcal{S}\subset\mathbb{R}.$ (1)
The most commonly used scoring rule is the estimated conditional average treatment effect (CATE) given the pre-treatment covariates, i.e., $f(\bm{x})=\widehat{\mathbb{E}}(Y_{i}(1)-Y_{i}(0)\mid\bm{X}_{i}=\bm{x})$, using a machine learning (ML) algorithm. Another common scoring rule is the risk score, which is typically represented by the baseline conditional outcome, i.e., $f(\bm{x})=\widehat{\mathbb{E}}(Y_{i}(0)\mid\bm{X}_{i}=\bm{x})$.
Until Section 2.7, we assume that a scoring rule is given. In some applications, the available scoring rule is trained on auxiliary data and is only an estimate of an idealized population scoring rule. One might, therefore, argue that inference about selected subgroup should integrate the sampling variability of the estimated scoring rule. Our goal, however, is to evaluate the empirical performance of a fixed scoring rule that can actually be deployed in the real world rather than the “true” scoring rule that can never be attained. For example, regulatory guidance for confirmatory trials emphasizes the prespecification of design choices, including any algorithms and associated cutoffs (Lewis, 1999). Similarly, the evaluation of machine learning tools in public policy, such as pretrial risk assessment instruments, calls for the assessment of actual recommendation system itself (Imai et al., 2023).
Thus, we first develop inference conditional on a fixed scoring rule, focusing on evaluation uncertainty rather than estimation uncertainty of the scoring rule. In Section 2.7, we discuss the condition, under which our inference remains valid even when one is interested in evaluating the empirical performance of the true population scoring rule. For now, the only assumption we impose is that the resulting treatment prioritization score is continuous.
Assumption 3 (Continuous Treatment Prioritization Score)
The cumulative distribution of the treatment prioritization score $F(s)=\Pr(S_{i}\leq s)$ is continuous.
In practice, we can break a tie, if it exists, by adding a small amount of noise to the score though our framework does not account for the uncertainty that arises from this random tie-breaking procedure.
Next, we make two regularity assumptions. First, we assume that the CATE, i.e., $\mathbb{E}(\psi_{i}\mid\bm{X}_{i})$ where $\psi_{i}=Y_{i}(1)-Y_{i}(0)$, is continuous in the quantile of the treatment prioritization score:
Assumption 4 (Continuity of the conditional average treatment effect)
The conditional average treatment effect given the quantile of the treatment prioritization score $\mathbb{E}[\psi_{i}\mid F(S_{i})=p]$ is continuous for $p\in[0,1]$.
A similar assumption is commonly invoked in the causal inference literature (e.g. Wager and Athey, 2018; Yadlowsky et al., 2021). In addition, we assume that the individual treatment effect has a finite second moment. The assumption is necessary for our asymptotic convergence result to have a well defined distribution.
Assumption 5 (Moment Condition for the Individual Treatment Effect)
We assume that the individual treatment effect $\psi_{i}$ has a finite second moment, i.e., $\mathbb{E}(\psi_{i}^{2})<\infty$.
2.2 Subgroup Identification with a Statistical Performance Guarantee
We consider the setting where a researcher wishes to utilize a treatment prioritization score to identify a subpopulation of individuals who benefit from treatment the most. The same methodology can be applied (by simply flipping the treatment prioritization score via $S_{i}=-f(\bm{X}_{i})$) to the case where researchers wish to identify a subgroup of individuals who are harmed most by treatment.
Formally, we wish to find the quantile $p^{\ast}$ of the true treatment prioritization score such that the sorted group average treatment effect (GATES) is maximized subject to a constraint set $\mathcal{P}\subseteq[0,1]$:
$p^{\ast}=\operatornamewithlimits{argmax}_{p\in\mathcal{P}}\Psi(p)\quad\text{where}\ \Psi(p)\ =\ \mathbb{E}[\psi_{i}\mid F(S_{i})\geq 1-p],$ (2)
where, under Assumption 2, the GATES is identified as,
$\Psi(p)=\mathbb{E}[Y_{i}\mid T_{i}=1,F(S_{i})\geq 1-p]-\mathbb{E}[Y_{i}\mid T_{i}=0,F(S_{i})\geq 1-p].$ (3)
If $\Psi(p)$ is known exactly, the optimization problem in Equation (2) leads to the selection of individuals, based on the quantiles of prioritization score, whose average treatment effect is the greatest. In practice, however, we only have a data set of finite sample size, and both $S_{i}$ and $\Psi(p)$ may be estimated with bias and noise.
Researchers often choose to solve the following empirical version of this optimization problem,
$\hat{p}_{n}\ =\ \operatornamewithlimits{argmax}_{p\in\mathcal{P}}\widehat{\Psi}_{n}(p),$ (4)
where
$\widehat{\Psi}_{n}(p)=\frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]}\quad\text{and }\ \hat{\psi}_{[n,i]}=\frac{T_{[n,i]}Y_{[n,i]}}{n_{1}/n}-\frac{(1-T_{[n,i]})Y_{[n,i]}}{n_{0}/n}$ (5)
with the subscript $[n,i]$ denoting the $i$th order statistic where the ordering is based on the treatment prioritization score $\hat{S}_{i}=\hat{f}(\bm{X}_{i})$. Thus, for example, if we use $\hat{S}_{[n,i]}$ to denote the $i$th order statistic of $\hat{S}$, then we have $\hat{S}_{[n,1]}\geq\hat{S}_{[n,2]},\ldots,\geq\hat{S}_{[n,n]}$.
While we can easily compute $\hat{p}_{n}$, the resulting estimate does not take into account the estimation uncertainty of $\widehat{\Psi}_{n}$. This is problematic especially when the sample size is limited and/or the true proportion of exceptional responders $p^{\ast}$ is small. In addition, because the selection of the parameter $p$ depends on the same data, the standard confidence band for $\widehat{\Psi}_{n}$ computed by conditioning on $\hat{p}_{n}$ suffers from a multiple testing problem and does not have a proper coverage guarantee.
To address these challenges, we consider the identification of exceptional responders with a statistical performance guarantee. We first develop an asymptotic uniform one-sided confidence band $C_{n}(p,\alpha)$ that covers the true value $\Psi(p)$ for any value of $p$ with the probability of at least $1-\alpha$, i.e.,
$\mathbb{P}\left(\forall p\in[0,1],\ \Psi(p)\geq\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha)\right)\geq 1-\alpha.$ (6)
We then use this uniform confidence band to obtain a statistical guarantee when selecting those who are predicted to benefit most from the treatment. Specifically, instead of selecting the value of $p$ that maximizes the estimated GATES as done in Equation (4), we choose the value of $p$ such that the lower confidence bound of the estimated GATES is maximized,
$\underline{\hat{p}}_{n}\ =\ \operatornamewithlimits{argmax}_{p\in[0,1]}\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha).$ (7)
This procedure gives a statistical guarantee that, with probability at least $1-\alpha$, the true GATES among those who are identified as benefiting most from the treatment is at least as great as the optimized lower confidence bound,
$\mathbb{P}\left(\Psi(p^{\ast})\geq\widehat{\Psi}_{n}(\underline{\hat{p}}_{n})-C_{n}(\underline{\hat{p}}_{n},\alpha)\right)\geq\mathbb{P}\left(\Psi(\underline{\hat{p}}_{n})\geq\widehat{\Psi}_{n}(\underline{\hat{p}}_{n})-C_{n}(\underline{\hat{p}}_{n},\alpha)\right)\geq 1-\alpha.$
where the first inequality follows from the fact that $\Psi(p^{\ast})\geq\Psi(p)$ for any $p$. Note that since we can never attain $p^{\ast}$, $\Psi(\underline{\hat{p}}_{n})$ is of more immediate interest as a target quantity for practitioners than $\Psi(p^{\ast})$.
Since this statistical guarantee does not depend on how the value of $p$ is selected, we can accommodate a wide variety of ways in which this threshold parameter is selected. For example, one may wish to identify the largest group of individuals whose GATES is at least as great as the pre-specified threshold $c$, i.e., $p^{\ast}(c)\ =\ \sup\{p\in[0,1]:\Psi(p)\geq c\}$. Suppose that we choose the value of $p$ such that the lower confidence bound matches with the threshold $c$, i.e.,
$\underline{\hat{p}}_{n}(c)\ =\ \sup\{p\in[0,1]:\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha)\geq c\}.$
Then, this procedure provides a statistical guarantee that, with the probability at least $1-\alpha$, the GATES for the selected group of individuals is at least as great as the threshold $c$.
$\mathbb{P}\left(\Psi(\underline{\hat{p}}_{n}(c))\geq\widehat{\Psi}_{n}(\underline{\hat{p}}_{n}(c))-C_{n}(\underline{\hat{p}}_{n},\alpha)\geq c\right)\geq 1-\alpha.$
Next, we show how to construct the uniform confidence band given in Equation (6).
2.3 Constructing Uniform Confidence Bands
To develop the uniform convergence result, we begin by showing that a normalized version of $p\widehat{\Psi}_{n}(p)$ converges in distribution to a zero-mean Gaussian process. In our proof, we utilize a generalized version of Donsker’s invariance principle for correlated random variables. For completeness, we state this theorem in Appendix Theorem S1, and refer readers to Borovkov and Sahanenko (1983) for its proof in the independent case and Ossiander (1987) for the correlated case. We now state the convergence result.
Proposition 1 (Convergence of the Normalized GATES Estimator)
Consider the random continuous polygon $s_{n}$ connected by the following basic points $b_{i}$ for $i=1,\cdots,n$:
$b_{i}=\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right).$
Under Assumptions 15, $s_{n}$ converges in distribution to a zero-mean Gaussian process uniformly in the Skorokhod space.
Proof is given in Appendix S1. Proposition 1 tells us that a properly normalized version of the GATES estimator does converge to an Gaussian Process - however it can have a complex covariance function that makes the calculation of uniform confidence bands difficult. Let us consider the correlation between $\widehat{\Psi}_{n}(p)$ and $\widehat{\Psi}_{n}(q)$ for any $p>q$. After normalizing them by their group size, we can write the difference between these two GATES estimators as,
$p\widehat{\Psi}_{n}(p)-q\widehat{\Psi}_{n}(q)\ =\ \frac{1}{n}\sum_{i=\lfloor nq\rfloor}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]}.$
where $\hat{\psi}_{[n,i]}$ is defined in Equation (5). Therefore, the correlation between $\widehat{\Psi}_{n}(p)$ and $\widehat{\Psi}_{n}(q)$ is determined by the correlation among $\hat{\psi}_{[n,i]}$. If the estimated treatment prioritization score $\hat{S}_{i}=\hat{f}(\bm{X}_{i})$ exactly equals the CATE, then $\psi_{[n,i]}$ will be strongly positively correlated with one another because they are a noisy version of the true CATE.
In particular, this dependence makes our problem distinct from the one studied in the previous literature, which focuses on testing whether an ordering $\hat{S}$ is correlated with the CATE (Miller and Siegmund, 1982; Hothorn and Lausen, 2002). In the hypothesis testing problem, the null hypothesis assumes that the $\psi_{[n,i]}$ are independent. This makes it possible to derive an asymptotically valid test statistic that follows the standard Weiner process. In contrast, since our goal is the construction of the uniform confidence intervals of the GATES, we impose no assumption about the dependence structure between $\hat{S}$ and CATE.
Somewhat surprisingly, we show that regardless of the accuracy of this score, the sorted individual treatment effect estimates $\hat{\psi}_{[n,i]}$ are non-negatively correlated with one another.
Proposition 2
(Non-negative Correlation of Sorted Individual Treatment Effect Estimates) Under Assumptions 15 and for all $1\leq i<j\leq n$, we have:
$\operatorname{Cov}(\hat{\psi}_{[n,i]},\hat{\psi}_{[n,j]})\geq 0.$
Proof is given in Appendix S2. The above inequality is tight and reaches its bound of 0 when the score $S_{i}$ is independent of the individual treatment effect $\psi_{i}$, i.e., when the ML estimate of the CATE is completely uninformative. We emphasize that this inequality is general and applicable to arbitrary evaluation metrics beyond GATES. This non-negative correlation is induced by the nature of the ordering operation.
Figure 1: Simulation study illustrating a non-negative correlation between the estimated treatment prioritization score $\hat{S}_{i}$ and the individual treatment effect $\psi_{i}$. Even when the score is negatively correlated with the individual treatment effect, the individual treatment effects ordered by the score are positively correlated. The magnitude of positive correlation increases as the relationship between the score and individual treatment effect becomes stronger. The correlation becomes zero when the score is independent of the individual treatment effect.
To illustrate this property, we simulate the data using $n=100$ and $Y_{i}(t)=X_{i}(1+t)$ for $t=0,1$, where $X_{i}\sim N(0,1)$, implying $\psi_{i}=X_{i}$. Then, we create the treatment prioritization score using $\hat{S}_{i}=rX_{i}+\sqrt{1-r^{2}}Z_{i}$ where $Z_{i}\sim N(0,1)$ so that $\operatorname{Corr}(\hat{S}_{i},\psi_{i})=r$ can be varied in $|r|\leq 1$. We estimate $\operatorname{Cov}(\hat{\psi}_{[n,i]},\hat{\psi}_{[n,i+1]})$ based on 10,000 Monte Carlo trials by averaging it across all $i$. Figure 1 illustrates the non-negative correlation property shown in Proposition 2 by plotting the estimated covariance between $\psi_{[n,i]}$ and $\psi_{[n,i+1]})$ against the estimated correlation between $S_{i}$ and $\psi_{i}$ after averaging across all $i$. As the correlation between the treatment prioritization score $S_{i}$ and the individual treatment effect becomes stronger, the magnitude of positive covariance between the sorted individual treatment effects increases. Even when the score is negatively correlated with the individual treatment effect, the correlation of sorted individual treatment effects remains positive.
Proposition 2 implies that a properly normalized version of $p\widehat{\Psi}_{n}(p)$ converges to a stochastic process with non-negatively correlated increments. This key insight enables us to bound the process with an appropriately rescaled Wiener process, which has independent increments. Specifically, we use Slepian’s lemma, which states that mean-zero Gaussian stochastic processes are more tightly concentrated when the increments have a greater covariance (see Lemma S3 in Appendix S3). This leads to the following construction of a uniform confidence band.
Theorem 1 (Uniform Minimum-area Confidence Band for the GATES Estimator)
Under Assumptions 15 and for a given level of significance $\alpha\in(0,1)$, we have,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\widehat{\Psi}_{n}(p)-\frac{\beta_{0}^{\ast}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}-\beta_{1}^{\ast}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(p))}\right)\geq 1-\alpha.$
where
$\mathbb{V}(\widehat{\Psi}_{n}(p))\ =\ \frac{1}{p^{2}}\left\{\frac{\mathbb{E}(S_{p1}^{2})}{n_{1}}+\frac{\mathbb{E}(S_{p0}^{2})}{n_{0}}-\frac{p(1-p)}{n-1}\mathbb{E}(\psi_{i}\mid\widehat{F}(\hat{S}_{i})\geq p)^{2}\right\},$
$S_{pt}^{2}=\sum_{i=1}^{n}(\hat{\phi}_{ip}(t)-\overline{\hat{\phi}_{p}(t)})^{2}/(n-1)$ with $\hat{\phi}_{ip}(t)=\mathbf{1}\{\widehat{F}(\hat{S}_{i})\geq p\}Y_{i}(t)$, and $\overline{\hat{\phi}_{p}(t)}=\sum_{i=1}^{n}\hat{\phi}_{ip}(t)/n$. Furthermore, we have:
$\{\beta_{0}^{\ast}(\alpha),\beta_{1}^{\ast}(\alpha)\}=\operatorname*{argmin}_{\beta_{0},\beta_{1}\in\mathbb{R}_{+}^{2}}\int^{1}_{0}\beta_{0}+\beta_{1}\sqrt{t}\ \mathrm{d}t\ \text{ subject to }\ \mathbb{P}\left(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\forall t\in[0,1]\right)\geq 1-\alpha.$ (8)
with $W(t)$ denotes the standard Wiener process.
Proof is given in Appendix S3. Similar to Proposition 2, the inequality is tight when the treatment prioritization score is completely uninformative about the CATE.
To solve Equation (8), note that the objective function is linear in $\beta_{0}$ and $\beta_{1}$ and the probability expression, i.e., $\mathbb{P}\left(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\forall t\in[0,1]\right)$, is monotonically increasing in these coefficients. Therefore, a modified binary search algorithm on $(\beta_{0},\beta_{1})$ on a box $(\beta_{0},\beta_{1})\in[0,\overline{\beta}_{0}]\times[0,\frac{3}{4}\overline{\beta}_{0}]$ suffices, where $\overline{\beta}_{0}=\operatorname*{argmin}_{\beta_{0}\in\mathbb{R}}\mathbb{P}(W(t)\leq\beta_{0}))\geq 1-\alpha$, and $\overline{\beta}_{0}$ can be found through a standard line search. An example algorithm that can be used to solve Equation (8) is presented as Algorithm 1 where we approximate the probabilities through Monte Carlo simulation.
Input: Upper bound for $\beta_{0}$, denoted by $\overline{\beta}_{0}$, step size $\epsilon$, Type I error probability $\alpha$
   Output: Optimal values of $\{\beta_{0}(\alpha),\beta_{1}(\alpha)\}$
[…]
2:for  $\beta_{0c}$ from 0 to $\overline{\beta}_{0}$ with step $\epsilon$ do
3:  $\beta_{1l}\leftarrow 0$
4:  while $\beta_{1h}-\beta_{1l}>\epsilon$ do
5:   $\beta_{1m}\leftarrow\frac{\beta_{1l}+\beta_{1h}}{2}$
6:   if $\mathbb{P}\left(W(t)\leq\beta_{0c}+\beta_{1m}\sqrt{t},\;\forall t\in[0,1]\right)>1-\alpha$ then
7:     $\beta_{1h}\leftarrow\beta_{1m}$
8:   else
9:     $\beta_{1l}\leftarrow\beta_{1m}$
10:   end if
11:  end while
12:  if $\int^{1}_{0}\beta_{0c}+\beta_{1m}\sqrt{t}\ \mathrm{d}t<V$ then
13:   $V\leftarrow\int^{1}_{0}\beta_{0c}+\sqrt{t}\beta_{1m}\ \mathrm{d}t$
14:   $\{\beta_{0}(\alpha),\beta_{1}(\alpha)\}\leftarrow\{\beta_{0c},\beta_{1m}\}$
15:  end if
16:end for
17:return $\{\beta_{0}(\alpha),\beta_{1}(\alpha)\}$
Algorithm 1 Bounding the Weiner process by minimizing the area.
The uniform confidence band given in Theorem 1 minimizes the area of the resulting confidence region by approximating it via a specific functional form $\beta_{0}+\sqrt{t}\beta_{1}$. This choice is inspired by Kendall et al. (2007) who find it to perform well for Brownian motion. We also consider a slight modification of this minimum-area confidence bands. Policymakers are rarely interested in extremely small $p$, as the sample size is too small to provide a conclusive evidence. Thus, we may only wish to minimize the area of confidence region over $p\in[p_{l},1]$ for some $p_{l}\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\hat{\Psi}_{n}(p)-\frac{\beta_{0}^{\ast}(\alpha;p_{l})}{p}\sqrt{\mathbb{V}(\hat{\Psi}_{n}(1))}-\beta_{1}^{\ast}(\alpha;p_{l})\sqrt{\mathbb{V}(\hat{\Psi}_{n}(p))}\right)\geq 1-\alpha,$
where
$\{\beta_{0}^{\ast}(\alpha;p_{l}),\beta_{1}^{\ast}(\alpha;p_{l})\}=\operatorname*{argmin}_{\beta_{0},\beta_{1}\in\mathbb{R}_{+}^{2}}\left\{\int^{1}_{p_{l}}\beta_{0}+\beta_{1}\sqrt{t}\ \mathrm{d}t\;\;\text{subject to}\;\;\mathbb{P}(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\;\forall t\in[0,1])\geq 1-\alpha\right\}.$ (9)
Figure 2: Two functional forms (Equations (9) and (10)) used to create an one-sided 95% uniform confidence band for Brownian motion. 20 sample paths of Brownian motion are plotted in gray for reference.
The above functional form is not unique, and many other valid choices of confidence bands exist. As an example, we next extend Theorem 1 to provide confidence bands using a function of the form $t^{k}$ for any $0\leq k<1/2$. Figure 2 illustrates these two functional forms used for the confidence bands of Brownian motion.
Corollary 1
Under Assumptions 15, we have, for a given level of significance $\alpha$, and any $0\leq k<1/2$,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\widehat{\Psi}_{n}(p)-\gamma^{\ast}(\alpha;k)\mathbb{V}(\widehat{\Psi}_{n}(p))^{k}\mathbb{V}(\widehat{\Psi}_{n}(1))^{1/2-k}p^{2k-1}\right)\geq 1-\alpha.$
where
$\gamma^{\ast}(\alpha;k)=\inf\{\gamma\in\mathbb{R}\mid\mathbb{P}(W(t)\leq\gamma t^{k},\;\forall\;t\in[0,1])\geq 1-\alpha\}.$ (10)
Proof is given in Appendix S4.
Note that $k$ can only take values less than $1/2$. This is because the law of iterated logarithm states,
$\limsup_{t\to 0}\frac{W(t)}{\sqrt{2t\log\log\frac{1}{t}}}\to 1$
which implies that, for any $\gamma>0$, $\mathbb{P}(W(t)\leq\gamma\sqrt{t},\;\forall\;t\in[0,1])=0$. Thus, $\gamma^{\ast}(\alpha;k)$ does not exist for any $k\geq 1/2$. Since heuristically we expect $\mathbb{V}(\hat{\Psi}_{n}(p))$ to scale $O\left(p^{-1}\right)$, the width of the confidence bands for a given $k$ should scale roughly as $O\left(p^{k-1}\right)$. Therefore, if a policymaker is interested in selecting exceptional responders for a relatively small value of $p$, one should choose $k$ as close to $1/2$ as possible.
2.4 Numerical Illustration
We present a numerical example to illustrate the uniform confidence bands of different forms introduced above. We utilize the data generating process described in Section 3 and train the LASSO regression model by fitting it to the population data based on the empirical distribution of the full sample. Then, we generate a random sample of size $n=2,500$ from this population and apply our methodology to the CATE estimates from the LASSO regression. Specifically, we compute the proposed minimum-area uniform confidence bands in two ways, based on Equation (9) with $p_{l}=0$ and $p_{l}=0.05$, and Equation (10) with $k=0$ and $k=0.499$, using 10,000 Monte Carlo trials to approximate the probabilities. We also provide the pointwise confidence band as a reference.
(a) Entire range $p\in[0,1]$
(b) Around $p=0.15$
(c) Around $p=0.65$
Figure 3: Comparison of uniform minimum-area confidence bands using the sample of size $n=2500$.
Figure 3(a) shows the resulting confidence bands with the solid black line representing the point estimate $\widehat{\Psi}_{n}$, whereas Figures 3(b) and 3(c) zoom into the areas around $p=0.15$ and $p=0.65$, respectively. We find that in this numerical example, the confidence band based on Equation (10) with $k=0$ (solid orange line) has the greatest width when $p$ is small although it is slightly tighter than the other uniform minimum-area confidence bands when $p$ is large. The other three confidence bands are quite similar although the one based on Equation (10) with $k=0.499$ is slightly wider (approximately 5%) than those based on Equation (9).
We also observe that in general, the different types of uniform confidence bands are on average approximately $50\%$ wider than the pointwise confidence bands. This suggests that the uniform bands can be practically useful while retaining a strong statistical guarantee. The uniform bands, relative to the pointwise ones, are generally wider when $p$ becomes closer zero, and tighter when $p$ gets closer to one. This is because the variability of $\widehat{\Psi}(p)$ decreases as $p$ increases.
2.5 Extension to Other Evaluation Metrics
So far, we have focused on the identification of exceptional responders that maximizes $\Psi(p)=\mathbb{E}[\psi_{i}\mid F(S_{i})\geq p]$, using the individual-level evaluation metric, $\hat{\psi}_{i}=T_{i}Y_{i}/(n_{1}/n)-(1-T_{i})Y_{i}/(n_{0}/n)$. We emphasize, however, that Theorem 1 is applicable to any unit-level statistic,
$\hat{\psi}_{i}^{g}\ =\ g(Y_{i},T_{i},\bm{X}_{i},p),$
so long as the continuity and finite moment conditions (i.e., Assumptions 4 and 5) hold. This generalization allows for various alternative formulations of exceptional responder problem. For example, we may adjust for the cost of treatment and use the following alternative definition of $\Psi$,
$\Psi^{C}(p)\ =\ \mathbb{E}[Y_{i}(1)-Y_{i}(0)-c_{p}(\bm{X}_{i})\mid F(S_{i})\geq 1-p]$
where $c_{p}(\cdot)$ is a known cost function based on the individual characteristics $\bm{X}_{i}$ and the proportion of exceptional responders $p$. Since this still leads to the individual level statistic $\hat{\psi}_{i}^{C}=\hat{\psi}_{i}-c_{p}(\bm{X}_{i})$, Theorem 1 is directly applicable.
Next, we consider the mean adjusted individual-level statistic, which can be written as,
$\hat{\psi}_{i}^{M}(\{Y_{i},T_{i},\bm{X}_{i}\}_{i=1}^{n},p)\ =\ h(Y_{i},T_{i},\bm{X}_{i},p)-\frac{p}{n}\sum_{i^{\prime}=1}^{n}h(Y_{i^{\prime}},T_{i^{\prime}},\bm{X}_{i^{\prime}},p),$
where $h(Y_{i},T_{i},\bm{X}_{i},p)$ is a unit-level statistic. Examples of such mean-adjusted statistics include the rank average treatment effect (RATE) (Yadlowsky et al., 2021) and the population average prescriptive effect (PAPE) (Imai and Li, 2023).
Here, motivated by the idea of the PAPE, we consider the following objective that adjusts for the overall average treatment,
$p^{\ast}\ =\ \operatornamewithlimits{argmax}_{p\in[0,1]}\Psi^{M}(p)\quad\text{where }\ \Psi^{M}(p)\ =\ \mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid F(S_{i})\geq 1-p]-p\mathbb{E}[Y_{i}(1)-Y_{i}(0)].$
Then, the unit-level statistic is given by,
$\hat{\psi}_{i}^{M}(\{Y_{i},T_{i},\bm{X}_{i}\}_{i=1}^{n},p)\ =\ \frac{Y_{i}T_{i}}{n_{1}/n}-\frac{Y_{i}(1-T_{i})}{n_{0}/n}-\left(\frac{p}{n_{1}}\sum_{i=1}^{n}Y_{i}T_{i}-\frac{p}{n_{0}}\sum_{i=1}^{n}Y_{i}(1-T_{i})\right).$
The sample estimator of $\Psi^{M}(p)$ can be written as follows,
$\widehat{\Psi}^{M}_{n}(p)\ =\ \frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}^{M}_{[n,i]}\ =\ \frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]}-\frac{p\lfloor np\rfloor}{np}\frac{1}{n}\sum_{i=1}^{n}\hat{\psi}_{[n,i]}=\widehat{\Psi}_{n}(p)-\frac{\lfloor np\rfloor}{n}\widehat{\Psi}_{n}(1).$ (11)
This alternative expression shows that $\widehat{\Psi}^{M}_{n}(p)$ can be constructed from $\widehat{\Psi}_{n}(p)$ in the same fashion as how a Brownian bridge on $[0,1]$ is constructed from a Wiener process. Therefore, Theorem 1 can be extended to the mean-adjusted statistic by replacing the Wiener process with the Brownian bridge.
Proposition 3
Under Assumptions 15, we have, for a given level of significance $\alpha\in(0,1)$,
$\lim_{n\to\infty}\mathbb{P}\left(\forall p\in[0,1],\Psi(p)\geq\widehat{\Psi}^{M}_{n}(p)-\frac{\delta_{0}^{\ast}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}-\delta_{1}^{\ast}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(p))\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)}\right)\geq 1-\alpha,$
where
$\{\delta_{0}^{\ast}(\alpha),\delta_{1}^{\ast}(\alpha)\}=\operatorname*{argmin}_{\delta_{0},\delta_{1}\in\mathbb{R}}\left\{\int^{1}_{0}\delta_{0}+\delta_{1}\sqrt{t(1-t)}\ \mathrm{d}t\ \;\;\text{s.t.}\;\;\ \mathbb{P}\left(B(t)\leq\delta_{0}+\delta_{1}\sqrt{t(1-t)},\;\forall\;t\in[0,1]\right)\geq 1-\alpha\right\},$
with $B(t)$ denoting a standard Brownian Bridge on $[0,1]$.
Appendix S5 presents a proof of this proportion, which exploits the fact that Brownian Bridge $B(t)$ must be 0 at both $t=0$ and $t=1$. Here, we cannot use Slepian’s lemma as in Theorem 1 to control the correlation between $\psi^{M}_{[n,i]}$, since after the mean adjustment we no longer have $\operatorname{Cov}(\psi^{M}_{[n,i]},\psi^{M}_{[n,j]})\geq 0$. Therefore, to control the covariance structure of the limiting Brownian Bridge, we need to prove a conditional version of Slepian’s inequality using Kahane’s inequality (Kahane, 1985).
2.6 Characterizing the Selected Subgroup
In practice, it is important to characterize the selected subgroup so that one can understand what covariate profiles are associated with exceptional responders. There are two strategies that can be used to facilitate this process while leveraging the proposed methodology. First, when estimating a scoring rule, we can restrict it to a class of interpretable rules, such as decision trees and linear models. Since our methodology does not impose any restriction on how scoring rules are estimated (they do not have to be even consistent), any interpretable function class can be used to improve the interpretability of scoring rule. Once an optimal scoring rule is estimated within this restricted class, one can directly apply the proposed methodology to obtain a uniform confidence interval for the resulting GATES.
Another possibility is to characterize the selected group using covariates (Chernozhukov et al., 2019). For example, we can estimate the following average value of covariate $X_{j}$ among the selected group,
$\mathbb{E}(X_{j}\mid F(S)\geq 1-p^{*}).$
Thus, we can treat $X_{j}$ as an auxiliary outcome. Define,
$\Psi^{X_{j}}(p)=\mathbb{E}(X_{j}\mid F(S)\geq 1-p),\quad\text{and}\quad\widehat{\Psi}^{X_{j}}_{n}(p)=\frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}X_{j,[n,i]},$
where $X_{j,[n,i]}$ denotes the $i$th order statistic of $X_{j}$ based on the estimated scoring rule $\hat{S}_{i}=\hat{f}(\bm{X}_{i})$. We then apply the Bonferroni correction to construct the jointly uniform confidence lower bound as,
$\mathbb{P}(\forall p\in[0,1],\Psi^{X_{j}}(p)\geq\widehat{\Psi}^{X_{j}}_{n}(p)-C^{X_{j}}_{n}(p,\alpha),\Psi(p)\geq\widehat{\Psi}_{n}(p)-C_{n}(p,\alpha))\geq 1-2\alpha$
where $C^{X_{j}}_{n}(p,\alpha)$ is defined analogously to $C_{n}(p,\alpha)$. The uniform confidence upper bound can be derived similarly.
2.7 Estimation Uncertainty of Scoring Rule
In this paper, we assume that the scoring rule is fixed and do not account for the estimation error. As explained in Section 2.1, this is because, in practice, one would like to evaluate the empirical performance of an actual scoring rule that will be deployed in real-world applications. In other words, we are not interested in the empirical performance of an idealized “true” scoring rule $f(x)$, to which an estimated scoring rule $\hat{f}(x)$ converges. Indeed, such a rule is never attainable and hence can never be deployed.
Nevertheless, if one is interested in evaluating the empirical performance of this hypothetical scoring rule, it is possible to use our uniform confidence bands by assuming both $f(x)$ and $\hat{f}(x)$ are sufficiently continuous, and the estimated prioritization score converges to the “true” prioritization score at a sufficiently fast rate.
Specifically, we replace Assumptions 3 and 4 with the following alternative assumptions.
Assumption 3A (Estimated prioritization score)
Define the cumulative distribution function (CDF) of the true treatment prioritization score as $F(s)=\Pr(S\leq s)$, and let $c(p)=F^{-1}(1-p)$ be the 1p1-pth percentile of the true treatment prioritization score. We assume that for all $p\in[0,1]$, there exist positive constants $t_{0}$, $c_{0}$, and $\kappa$ such that we have for all $t\in[0,t_{0})$:
$\mathbb{P}(|S-c(p)|<t)\leq c_{0}t^{\kappa}$
Furthermore, we assume that the error of the estimated treatment prioritization score $\hat{f}(x)$ is bounded as follows:
$\sup_{x\in\mathcal{X}}|\hat{f}(x)-f(x)|=o(n^{-1/2\kappa})$
Assumption 4A (Continuity of the conditional average treatment effect)
The conditional average treatment effect given the quantile of the true and estimated treatment prioritization scores, $\mathbb{E}[\psi_{i}\mid F(S_{i})=p]$ and $\mathbb{E}[\psi_{i}\mid\hat{F}(\hat{S}_{i})=p]$, are continuous for $p\in[0,1]$.
Assumption 4A is a straightforward extension of the condition in Assumption 4, Assumption 3A has two components; one is a uniform margin condition on the true treatment prioritization score, and the other is a rate condition on the convergence of the estimated prioritization score. The use of margin conditions is common in the literature (see e.g. Qian and Murphy, 2011; Kitagawa and Tetenov, 2018; Bonvini et al., 2023; Ben-Michael et al., 2024).
These two conditions have a tradeoff relationship. In general, as long as the density of the true treatment prioritization score is bounded, we can have $\kappa=1$. However, with $\kappa=1$, we require the parametric rate of convergence, $o(n^{-1/2})$, that may not be satisfied by modern machine learning algorithms with high-dimensional covariates (Kennedy, 2023). Alternatively, if we have a stronger margin condition with $\kappa>1$, then we can utilize machine learning algorithms that converge slower than $n^{-1/2}$. Note that in general, it is impossible to have $\kappa>1$ for all $p\in[0,1]$, as it would imply a vanishing density for the true treatment prioritization score. However, if the decision-maker knows that the final $p$ would fall into a subset $\mathcal{P}\subset[0,1]$, it is possible that we can have the margin condition hold for $\kappa>1$ for $p\in\mathcal{P}$.
Under these alternative assumptions, it is possible to prove the equivalent Donsker result that is needed to prove the validity of our uniform confidence bands with the estimated treatment prioritization score.
Proposition 4
(Convergence of the normalized GATES estimator with the estimated treatment prioritization score) Consider the random continuous polygon $s_{n}$ connected by the following basic points $b_{i}$ for $i=1,\cdots,n$:
$b_{i}=\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right).$
Under Assumptions 1, 2, 3A, 4A, and 5, $s_{n}$ converges in distribution to a zero-mean Gaussian process uniformly in the Skorokhod space.
The proof is given in Appendix S6.
3 A Simulation Study
We conduct a simulation study to examine the finite sample performance of the proposed methodology. We utilize one data generating process from the 2016 Atlantic Causal Inference Conference (ACIC) Competition. We briefly describe its simulation setting here and refer interested readers to Dorie et al. (2019) for additional details. The focus of this competition was the inference of the average treatment effect in observational studies. There are a total of 58 pre-treatment covariates, including 3 categorical, 5 binary, 27 count data, and 13 continuous variables. The data were taken from a real-world study with the sample size $n=4802$.
In our simulation, we assume that the empirical distribution of these covariates represent the population of interest and obtain each simulation sample via bootstrap. One important change from the original competition is that instead of utilizing a propensity model to determine $T$, we assume that the treatment assignment is completely randomized, i.e., $\Pr(T_{i}=1)=1/2$, and the treatment and control groups are of equal size, i.e., $n_{1}=n_{0}=n/2$.
To generate the outcome variable, we use one of the data generating processes used in the competition (the 28th), which is based on the generalized additive model with polynomial basis functions. The model represents a setting, in which there exists a substantial amount of treatment effect heterogeneity. For the sake of completeness, the formula for this outcome model is reproduced here:
$\mathbb{E}(Y_{i}(t)\mid\bm{X}_{i})$ $=1.60+0.53\times x_{29}-3.80\times x_{29}(x_{29}-0.98)(x_{29}+0.86)-0.32\times\mathbf{1}\{x_{17}>0\}$
$+0.21\times\mathbf{1}\{x_{42}>0\}-0.63\times x_{27}+4.68\times\mathbf{1}\{x_{27}<-0.61\}-0.39\times(x_{27}+0.91)\mathbf{1}\{x_{27}<-0.91\}$
$+0.75\times\mathbf{1}\{x_{30}\leq 0\}-1.22\times\mathbf{1}\{x_{54}\leq 0\}+0.11\times x_{37}\mathbf{1}\{x_{4}\leq 0\}-0.71\times\mathbf{1}\{x_{17}\leq 0,t=0\}$
$-1.82\times\mathbf{1}\{x_{42}\leq 0,t=1\}+0.28\times\mathbf{1}\{x_{30}\leq 0,t=0\}$
$+\{0.58\times x_{29}-9.42\times x_{29}(x_{29}-0.67)(x_{29}+0.34)\}\times\mathbf{1}\{t=1\}$
$+(0.44\times x_{27}-4.87\times\mathbf{1}\{x_{27}<-0.80\})\times\mathbf{1}\{t=0\}-2.54\times\mathbf{1}\{t=0,x_{54}\leq 0\}.$
for $t=0,1$ where the error term is distributed according to the standard normal distribution.
We estimate the CATE and use it as our treatment prioritization score to identify exceptional responders. For the estimation of the CATE, we use Bayesian Additive Regression Trees (BART) (see Chipman et al., 2010; Hill, 2011; Hahn et al., 2020) and Causal Forest (Wager and Athey, 2018), and LASSO (Tibshirani, 1996). The models are trained by fitting them to the population data (i.e., the empirical distribution of the full sample). For implementation, we use R 3.6.3 with the following packages: bartMachine (version 1.2.6) for BART, grf (version 2.0.1) for Causal Forest, and glmnet (version 4.1-2) for LASSO. The number of trees was tuned through 5-fold cross-validation for BART and Causal Forest. The regularization parameter was tuned similarly for LASSO.
[…]
$\bm{n=100}$ $\bm{n=500}$ $\bm{n=2500}$ $\bm{n=100}$ $\bm{n=500}$ $\bm{n=2500}$ $\bm{n=100}$ $\bm{n=500}$ $\bm{n=2500}$
BART $96.1\%$ $96.0\%$ $95.2\%$ $87.2\%$ $76.5\%$ $70.3\%$ $96.5\%$ $96.7\%$ $97.1\%$
Causal Forest $96.0\%$ $95.3\%$ $95.7\%$ $83.7\%$ $77.1\%$ $71.9\%$ $96.2\%$ $97.2\%$ $98.0\%$
LASSO $95.8\%$ $\% 95.695.6\%$ $84.1\%$ $76.0\%$ $69.8\%$ $96.5\%$ $97.4\%$ $97.0\%$
Table 1: Empirical coverage of 95% confidence bands. The table presents the empirical coverage of uniform 95% confidence bands as presented in Theorem 1, along with its comparison to the coverage of the 95% pointwise confidence bands and $50\%$ inflated pointwise confidence bands. Coverage is computed as the percentage of experiments in which the confidence bands contains the entire GATES curve.
We investigate the finite-sample performance of the uniform confidence bands as defined in Theorem 1. Table 1 shows the proportion of simulation experiments, in which the confidence band contains the entire curve of $\Psi^{\ast}(p)$. We find that our uniform confidence bands reach the nominal coverage for every dataset size down to $n=100$ while the pointwise confidence bands suffer from a significant undercoverage, especially as the sample size $n$ grows. Although our estimators are theoretically conservative, we see little evidence of significant over-coverage across all algorithms and data sizes.
Following our observations in Section 2.4, we also show the coverage of the pointwise confidence bands after inflating it by $50\%$. We observe that this simple linear adjustment to pointwise confidence bands appear to achieve adequate coverage, suggesting that this may serve a heuristic but reasonable approximation to the $95\%$ uniform confidence bands.
Figure 4: Comparison of the lengths of the 95% uniform and pointwise confidence bands. The treatment prioritization score in this simulation is constructed using LASSO. The horizontal axis is the quantile threshold value of the score.
Finally, Figure 4 compares the lengths of these uniform and pointwise confidence intervals where the treatment prioritization score is estimated using LASSO. In general, the uniform confidence interval is about $25-50\%$ wider than the pointwise confidence intervals across different sample sizes.
4 Empirical Application
We apply the proposed methodology to the data from a clinical trial of late-stage prostate cancer.
4.1 Setup
Byar and Green (1980) present the data from a randomized placebo-controlled clinical trial to estimate the effect of diethylstilbestrol on overall survival of patients. Three different dosage levels (0.2 mg, 1.0mg, or 5.0mg) were used in the trial that involved 502 patients with stage 3 or 4 prostate cancer. For each patient, the trial recorded their baseline health characteristics, common laboratory measurements, prior disease history, and detailed information on the current prostate cancer. The primary outcome is the number of months of total survival at the end of follow-up, which may have occurred at either death or the program completion.
We will focus on the 5.0mg subgroup and consider the binary treatment of giving either 5.0mg estrogen or placebo for these patients. In total, 127 patients received the placebo while 125 patients received the estrogen treatment. The average treatment effect was estimated to be 0.3-0.3 months but this estimate was not statistically significant, suggesting that on average the estrogen treatment do not improve total survival.
By analyzing this data, we wish to identify a subgroup that may significantly benefit from the treatment although the overall average treatment effect estimate was not statistically significant. We utilize a total of 13 pre-treatment covariates, including age, weight, height, blood pressure, tumor size and stage, ECG diagnosis, and blood measurements. The estrogen treatment is denoted as $T_{i}=1$ whereas the placebo is represented as $T_{i}=0$. The outcome variable $Y_{i}$ is the total survival.
We randomly selected approximately 40% of the sample (i.e., 103 observations) as the training data and the reminder of the sample (i.e., 149 observations) as the evaluation data. We train four machine learning models: the same three machine learning models as those used in Section 3 and the Causal Tree model (Athey and Imbens, 2016), which is included to illustrate how the proposed methodology can be used to select interpretable subgroups of exceptional responders.
For Causal Forest, we tuned the number of trees on a grid of $[500,700,900,1100,1300,1500]$. For BART, tuning was done on the number of trees on a grid of $[50,70,90,110,130,150]$. For LASSO, we tuned the regularization parameter on a grid of $[0.01,0.05,0.1,0.5,1,5]$. For the Causal Tree, we tuned the minimum node size parameter on a grid of $[5,10,20,50]$. All tuning was done with a 5-fold cross-validation procedure on the training set with mean-squared error as the validation metric.
4.2 Findings
Figure 5: Evaluation of different machine learning algorithms on the randomized clinical trial for Diethylstilbestrol. The dark blue solid line represents the point estimates of the number of months of the increase in total survival when maximally $p\%$ of the population is treated with Diethylstilbestrol. The dotted red line represents the pointwise $90\%$ confidence interval while the light blue dashed line represents the uniform $90\%$ confidence interval. The black horizontal line represents the average treatment effect, which is 0.3-0.3 months.
Figure 5 shows the results. For each plot, we compute the point estimates of the GATES (solid blue lines), the 90% minimum-area lower confidence bands (blue dotted lines) as shown in Equation (8), and 90% pointwise confidence bands (red dotted lines). We find that in this application, BART and especially Causal Forest appear to be able to identify some patterns of treatment effect heterogeneity in the population. In contrast, LASSO and Causal Tree are unable to capture any statistically significant heterogeneity. We also find that the minimum-area uniform confidence bands are close to the pointwise bands for a larger value of $p$ and only diverge significantly for a small value of $p$.
Estimated proportion of Estimated 90% uniform
Estimator exceptional responders GATES confidence band
Causal Forest $39.6\%$ $17.9$ (4.57, $\infty$)
BART $29.5\%$ $19.9$ (3.31, $\infty$)
LASSO $82.6\%$ $3.04$ (-5.17, $\infty$)
Causal Tree $54.3\%$ $7.38$ (-3.69, $\infty$)
Table 2: Estimated proportion of exceptional responders, their GATES estimates, and 90% uniform confidence bands. Each row represents the results based on different machine learning algorithms used to estimate the treatment prioritization score. The confidence intervals are corrected for multiple-testing using the Bonferroni method. The outcome is total survival in months.
Table 2 presents the estimated proportion of exceptional responders, their GATES estimates, and the corresponding 90% uniform confidence bands, for each ML algorithm used to construct the treatment prioritization score. The results show that both BART and Causal Forest appears to be able to identify a subset of population that would experience a statistically significant positive effect of the treatment on total survival. Specifically, we find that among these subsets of patients, 5mg of estrogen will increase their total survival by more than 15 months on average. BART in particular finds that $29\%$ of the population that would exhibit an increase of 20 months in total survival on average with the 90% uniform lower confidence band of 3.31 months.
In comparison, the previous best optimization approach (e.g. Bertsimas et al., 2019) found a similarly sized group that could prolong total survival by 18 months on average but without a statistical guarantee. Although the exact subgroup might vary over multiple train-test splits, this example demonstrates that our methodology is able to exploit the power of modern machine learning algorithms while having a rigorous performance guarantee.
Figure 6: The scoring function $\hat{S}$ according to the causal tree model. If the condition at the split is satisfied, the path continues left, and vice versa. The number represents the average treatment effect at the node.
In contrast, LASSO and Causal Tree are unable to identify any subgroup with a statistically significant positive treatment effect. However, a main advantage of Causal Tree is that it produces an interpretable decision rule. As shown in Figure 6, its scoring function is a simple function of covariates. Notably, in the selection of its exceptional responders (of size $54.3\%$ of the original group), the procedure includes nearly all patients in the two terminal nodes with the highest estimated average treatment effects. This yields a transparent description of the exceptional responder subgroup: patients with large tumors, and patients with low-grade tumors and no history of cardiovascular disease.
Covariates Average Difference One-sided 90% CI
Tumor Size (cm2) $4.2$ $(0.4,\infty)$
Stage/Grade Index $2.1$ $(-0.1,\infty)$
Weight 3.7-3.7 $(-\infty,6.5)$
Table 3: The difference in average values of covariates between the BART-selected subgroup and the overall population. We report their (one-sided) 90% uniform confidence bands. The confidence bands are corrected for multiple-testing using the Bonferroni method.
To interpret the results of the other models, we can follow the strategy described in Section 2.6 and characterize the average values of various features within the selected subgroup. Table 3 shows the average difference between selected subgroup and the overall population (using the mean-adjusted statistic) for three covariates with 90% uniform confidence bands. The results are corrected for multiple testing, and the BART results are used for illustration. We find that the average tumor size of the selected group is 4.2 cm2 larger than the population average, and this is statistically significant at the $\alpha=10\%$ level. Similarly, we find that the stage/grade index, a combination index of the tumor stage and histologic grade, is larger, and the weight of the subgroup is lower than the population, though both differences are not statistically significant. This procedure allows an interpretable characterization of the subgroup that could lead to better understanding of the mechanisms behind exceptional responders.
5 Concluding Remarks
As the use of machine learning tools become widespread in high-stake decision making, it is essential to deploy them effectively and safely. Doing so requires the accurate quantification of statistical uncertainty and the robust evaluation of their empirical performance. The proposed methodology provides a statistical performance guarantee when selecting an optimal subset of a target population who are predicted to benefit most from a treatment of interest. We believe that our methodology is particularly useful when the evaluation sample is of limited size, machine learning models provide biased or noisy estimates of treatment efficacy, and/or the proportion of exceptional responders is relatively small. In future research, we plan to extend the proposed methodology to other settings, including non-binary and dynamic treatments. It is also of interest to consider the question of how to account for the statistical uncertainty that arises from the estimation of treatment prioritization score, which is taken as given in our proposed framework.
Supplementary Appendix
Appendix S1 Proof of Proposition 1
We begin by proving two lemmas and stating Donsker’s Invariance Principle, which we use to prove Proposition 1.
Lemma S1
Under Assumptions 14 and for some $\epsilon>0$,
$\lim_{n\to\infty}\sup_{p\in[0,1]}n^{1/2+\epsilon}\left(\mathbb{E}[p\widehat{\Psi}_{n}(p)]-p\Psi(p)\right)=0.$
Proof.
We bound the bias of $\mathbb{E}(\widehat{\Psi}_{n}(p))$ by appealing to Theorem 1 of Imai and Li (2023), which implies,
$|p\mathbb{E}(\widehat{\Psi}_{n}(p))-p\Psi(p)|\leq$ $\left|\mathbb{E}\left[\int_{F(c(p))}^{F(\hat{c}_{n}(p))}\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(x))\mathrm{d}x\right]\right|$ (S1)
where $\hat{c}_{n}(p)\ =\ \inf\{c\in\mathbb{R}:\sum_{i=1}^{n}\mathbf{1}\{S_{i}<c\}\geq n(1-p)\}$ denote the estimated $p$-quantile and $c(p)=F^{-1}(1-p)$ is the true $p$-quantile. By the definition of $\hat{c}_{n}(p)$, $F(\hat{c}_{n}(p))$ is the n(1p)n(1-p)th order statistic of $n$ independent uniform random variables. Therefore, it follows the Beta distribution with the shape and scale parameters equal to n(1p)n(1-p) and n+1n(1p)n+1-n(1-p), respectively. Now, by Assumption 4, we can compute the first-order Taylor expansion of $\int^{x}_{a}\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(x))\mathrm{d}x$ and thus:
[…]
$\biggl{.}\biggl{.}\hskip 72.26999pt+o(F(\hat{c}_{n}(p))-F(c(p)))\biggr{]}\biggr{|}$
$=$ $\sup_{p\in[0,1]}\left|\mathbb{E}(\hat{\psi}_{i}\mid S_{i}=F^{-1}(1-p))\right|\left|\frac{n(1-p)}{n+1}-(1-p)\right|+o\left(\frac{1}{n}\right)$
$=$ $O\left(\frac{1}{n}\right).$
In particular, we thus have, for any $0<\epsilon<\frac{1}{2}$:
$\lim_{n\to\infty}n^{\frac{1}{2}+\epsilon}\sup_{p\in[0,1]}\left(\mathbb{E}[p\widehat{\Psi}_{n}(p)]-p\Psi(p)\right)=0.$
Lemma S2
Under Assumptions 14, we have for some $\epsilon>0$:
$\lim_{n\to\infty}n^{1/2+\epsilon}\sup_{p\in[0,1]}\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]\;\mathrm{d}x=0$
Proof.
By definition of the floor function, we have that $(1-p)-\frac{\lfloor n(1-p)\rfloor}{n}\leq\frac{1}{n}$. Therefore for some $\epsilon<\frac{1}{2}$:
$\lim_{n\to\infty}n^{1/2+\epsilon}\sup_{p\in[0,1]}\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]dx\ \leq\ \lim_{n\to\infty}n^{-1/2+\epsilon}\sup_{p\in[0,1]}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=1-p]$
By Assumption 4, we know that $\sup_{p\in[0,1]}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=p]$ is finite. Therefore, the right handside of the above equation is zero, implying the desired result. ∎
Theorem S1 (Donsker’s Invariance Principle)
Consider a double sequence of random variables $U_{i,j}$ with $\mathbb{E}[U_{i,j}]=0$ for all $i,j\in\{1,2,\ldots,n\}$. Assume that $U_{n,i}$ satisfies the metric entropy integrability condition in $L_{2}$ as defined in Equation (3.1) of Ossiander (1987). Then, define the quantity:
$S_{n}(t)=\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}U_{n,i}$
where $t\in[0,1]$ and $S_{n}(0)=0$. Construct the random continuous polygon $s_{n}$ with basic points:
$b_{i}\ =\ \left(\frac{\mathbb{V}(S_{n}(i/n))}{\mathbb{V}(S_{n}(1))},\frac{S_{n}(i/n)}{\sqrt{\mathbb{V}(S_{n}(1))}}\right),$
where $i\in\{1,\cdots,n\}$. Then, as the sample size $n$ goes to infinity, $s_{n}$ converges in distribution to $G$ uniformly in the Skorokhod space, where $G$ is a zero-mean Gaussian process with covariance function
$K(p,p^{\prime})=\lim_{n\to\infty}\operatorname{Cov}\left(\frac{S_{n}(\lfloor pn\rfloor/n)}{\sqrt{\mathbb{V}(S_{n}(\lfloor pn\rfloor/n))}},\frac{S_{n}(\lfloor p^{\prime}n\rfloor/n)}{\sqrt{\mathbb{V}(S_{n}(\lfloor p^{\prime}n\rfloor/n))}}\right),$
in particular $K(p,p)=p$. Additionally, when $\operatorname{Cov}(U_{n,i},U_{n,j})=0$ for all $n$ and $1\leq i,j\leq n$, then $G$ is the standard Wiener process.
Finally, using the above results, we prove Proposition 1. Define for all $i\in\{1,\cdots,n\}$:
$\Delta\Psi_{n,i}=\frac{i}{n}\mathbb{E}\left[\psi_{i}\ \biggl{|}\ F(S_{i})\geq\frac{i}{n}\right]-\frac{i-1}{n}\mathbb{E}\left[\psi_{i}\ \biggl{|}\ F(S_{i})\geq\frac{i-1}{n}\right].$
Note that:
$p\widehat{\Psi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]},\qquad p\Psi(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\Delta\Psi_{n,i}+\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]\;\mathrm{d}x$
By Lemmas S1 and S2, we can construct a new random variable $U_{n,i}$ with $\mathbb{E}(U_{n,i})=0$ for some fixed quantity $\epsilon_{n,i}>0$ in the following manner,
$U_{n,i}$ $=\Delta\Psi_{n,i}-\hat{\psi}_{[n,i]}-\epsilon_{n,i},\quad\text{where}\ \sup_{t\in[0,1]}\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\epsilon_{n,i}=o(n^{-1/2}).$ (S2)
By Assumption 5 $\psi_{n,i}$ have finite second moments, so the induced order statistics $\psi_{[n,i]}$ satisfy the metric entropy integrability condition on $L_{2}$ (Davydov and Egorov, 2000). Therefore we can apply Theorem S1 to $U_{n,i}$ by considering a random continuous polygon $\tilde{s}_{n}$ connected by the basic points:
$\tilde{b}_{i}\ =\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\hat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})-\frac{1}{n}\sum_{i^{\prime}=1}^{i}\epsilon_{n,i^{\prime}}}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right),$
implying that $\tilde{s}_{n}$ converges to a zero-mean Gaussian Process. Equation (S2) implies $\frac{\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\epsilon_{n,i}}{\sqrt{\mathbb{V}(\Psi(1))}}\to 0$ uniformly in $t$. Thus, the random continuous polygon $s_{n}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right),$
converges to a zero-mean Gaussian process $G$. ∎
Appendix S2 Proof of Proposition 2
Proof.
Define $\psi(s)=\mathbb{E}(\psi_{i}\mid\hat{S}_{i}=s)$. Then, we have,
$\operatorname{Cov}(\hat{\psi}_{[n,i]},\hat{\psi}_{[n,j]})$
[…]
where $\hat{S}_{[n,i]}$ is the $i$th order statistic of $\hat{S}_{i}$ given the sample of size $n$. We follow the proof strategy of (Cuadras, 2002, Theorem 1) by defining $(\hat{S}_{[n,i]}^{\prime},\hat{S}_{[n,j]}^{\prime})$ as random variables that are independently and identically distributed as $(\hat{S}_{[n,i]},\hat{S}_{[n,j]})$. Then we have:
$\operatorname{Cov}(\psi(\hat{S}_{[n,i]}),\psi(\hat{S}_{[n,j]}))$ $=\mathbb{E}[\psi(\hat{S}_{[n,i]})\psi(\hat{S}_{[n,j]})]-\mathbb{E}[\psi(\hat{S}_{[n,i]})]\mathbb{E}[\psi(\hat{S}_{[n,j]})]$
$=\frac{1}{2}\mathbb{E}[(\psi(\hat{S}_{[n,i]})-\psi(\hat{S}_{[n,i]}^{\prime}))(\psi(\hat{S}_{[n,j]})-\psi(\hat{S}_{[n,j]}^{\prime}))]$
Since $\int_{-\infty}^{\infty}\mathbbm{1}(a\leq\hat{S}_{[n,i]})-\mathbbm{1}(a\leq\hat{S}_{[n,i]}^{\prime})\mathrm{d}\psi(a)=\psi(\hat{S}_{[n,i]})-\psi(\hat{S}_{[n,i]}^{\prime})$, we have:
$\operatorname{Cov}(\psi(\hat{S}_{[n,i]}),\psi(\hat{S}_{[n,j]}))$
$=$ $\frac{1}{2}\mathbb{E}\left[\int_{R^{2}}\left\{\mathbbm{1}(a\leq\hat{S}_{[n,i]})-\mathbbm{1}(a\leq\hat{S}_{[n,i]}^{\prime})\right\}\left\{\mathbbm{1}(b\leq\hat{S}_{[n,j]})-\mathbbm{1}(b\leq\hat{S}_{[n,j]}^{\prime})\right\}\mathrm{d}\psi(a)\mathrm{d}\psi(b)\right].$
Under the assumptions, $\psi(\cdot)$ is a continuous function and $\psi(\hat{S}_{[n,i]})$ has a finite second moment. Therefore, we can apply Fubini’s theorem to exchange integration signs:
$\operatorname{Cov}(\psi(\hat{S}_{[n,i]}),\psi(\hat{S}_{[n,j]}))=$ $\int_{R^{2}}1-\mathbb{P}(\hat{S}_{[n,i]}<a)\mathbb{P}(\hat{S}_{[n,j]}<b)+\mathbb{P}(\hat{S}_{[n,i]}<a,\hat{S}_{[n,j]}<b)$
$\quad-(1-\mathbb{P}(\hat{S}_{[n,i]}<a))(1-\mathbb{P}(\hat{S}_{[n,j]}<b))\mathrm{d}\psi(a)\mathrm{d}\psi(b)$
$=$ $\int_{R^{2}}\mathbb{P}(\hat{S}_{[n,i]}<a,\hat{S}_{[n,j]}<b)-\mathbb{P}(\hat{S}_{[n,i]}<a)\mathbb{P}(\hat{S}_{[n,j]}<b)\mathrm{d}\psi(a)\mathrm{d}\psi(b).$
Since $\hat{S}_{[n,i]}$ and $\hat{S}_{[n,j]}$ are order statistics, by Theorem 3.4 of Boland et al. (1996), we know that $\mathbb{P}(\hat{S}_{[n,i]}<a,\hat{S}_{[n,j]}<b)\geq\mathbb{P}(\hat{S}_{[n,i]}<a)\mathbb{P}(\hat{S}_{[n,j]}<b)$ for all $a,b$. Thus, the desired conclusion holds. ∎
Appendix S3 Proof of Theorem 1
To prove Theorem 1, we first state Slepian’s lemma.
Lemma S3 (Slepian (1962))
For Gaussian random variables $X=(X_{1},\cdots,X_{n})$ and $Y=(Y_{1},\cdots,Y_{n})$ in $\mathbb{R}^{n}$ satisfying $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]=0$, $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]$ and $\mathbb{E}[X_{i}X_{j}]\leq\mathbb{E}[Y_{i}Y_{j}]$ for all $i,j=1,\cdots,n$, we have for any real numbers $u_{1},\cdots u_{n}$:
$\mathbb{P}\left[\bigcap_{i=1}^{n}\{X_{i}\geq u_{i}\}\right]\leq\mathbb{P}\left[\bigcap_{i=1}^{n}\{Y_{i}\geq u_{i}\}\right].$
We define a set of independent Gaussian variables $\phi_{[n,i]}$ whose mean and variance are identical to those of $\hat{\psi}_{[n,i]}$, i.e., $\mathbb{E}(\phi_{[n,i]})=\mathbb{E}(\hat{\psi}_{[n,i]})$, $\mathbb{V}(\phi_{[n,i]})=\mathbb{V}(\hat{\psi}_{[n,i]})$, and $\operatorname{Cov}(\phi_{[n,i]},\phi_{[n,j]})=0$. Then, define:
$p\widehat{\Phi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\phi_{[n,i]}.$
Since $\phi_{[n,i]}$ are independent, the metric entropy integrability condition on $L_{2}$ is automatically satisfied by Assumption 5. Applying Theorem S1, the random continuous polygon $s_{n}^{\prime}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Phi}_{n}(1))},\frac{\frac{i}{n}\Phi(\frac{i}{n})-\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Phi}_{n}(1))}}\right)$
converges to $W$, where $W$ is the standard Wiener process. Then, we have:
$\mathbb{P}\left(W(p)\geq-\beta_{0}^{\ast}(\alpha)-\beta_{1}^{\ast}(\alpha)\sqrt{p},\;\forall\;p\in[0,1]\right)\ =\ 1-\alpha,$ (S3)
where
$\{\beta_{0}^{\ast}(\alpha),\beta_{1}^{\ast}(\alpha)\}=\operatorname*{argmin}_{\beta_{0},\beta_{1}\in\mathbb{R}_{+}^{2}}\left\{\int^{1}_{0}\beta_{0}+\beta_{1}\sqrt{t}\ \mathrm{d}t\;\;\text{subject to}\;\;\mathbb{P}(W(t)\leq\beta_{0}+\beta_{1}\sqrt{t},\;\forall\;t\in[0,1])\geq 1-\alpha\right\}.$
Proposition 2 implies that the covariance function $K_{G}$ of the Gaussian Process $G$ satisfies $K_{G}(p,p^{\prime})\geq K_{W}(p,p^{\prime})=0$ where $K_{W}$ is the covariance function of $W$. In addition, we have $K_{G}(p,p)=K_{W}(p,p)$ by construction. Thus, we can apply Slepian’s lemma (Lemma S3) to Equation (S3) and obtain,
$\mathbb{P}\left(G(p)\geq-\beta^{*}_{0}(\alpha)-\beta^{*}_{1}(\alpha)\sqrt{p},\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$ (S4)
Proposition 1 implies that uniformly for all $p\in[0,1]$, we have as $n\to\infty$:
$\left(\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right)\to(p,G(p))$
Therefore, we can rewrite Equation (S4) as:
$\lim_{n\to\infty}\mathbb{P}\left(\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\geq-\beta^{*}_{0}(\alpha)-\sqrt{\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\beta^{*}_{1}(\alpha),\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$
Rearranging the terms gives:
$\lim_{n\to\infty}\mathbb{P}\left(p\Psi(p)\geq p\widehat{\Psi}_{n}(p)-\beta^{*}_{0}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}-\beta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(p\widehat{\Psi}_{n}(p))},\;\forall p\in[0,1]\right)\geq 1-\alpha.$
Therefore, for any $c_{0}>0$, we have:
$\lim_{n\to\infty}\mathbb{P}\left(\Psi(p)\geq\widehat{\Psi}_{n}(p)-\frac{\beta^{*}_{0}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}-\beta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(p))},\;\forall\;p\in[c_{0},1]\right)\geq 1-\alpha.$
Taking $c_{0}\to 0$ gives the required result. The expression of $\mathbb{V}(\widehat{\Psi}_{n}(p))$ follows directly from Theorem 1 of Imai and Li (2023). ∎
Appendix S4 Proof of Corollary 1
Identical to Theorem 1, we define a set of independent Gaussian variables $\phi_{[n,i]}$ whose mean and variance are identical to those of $\hat{\psi}_{[n,i]}$, i.e., $\mathbb{E}(\phi_{[n,i]})=\mathbb{E}(\hat{\psi}_{[n,i]})$, $\mathbb{V}(\phi_{[n,i]})=\mathbb{V}(\hat{\psi}_{[n,i]})$, and $\operatorname{Cov}(\phi_{[n,i]},\phi_{[n,j]})=0$. Then, define:
$p\widehat{\Phi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\phi_{[n,i]}.$
Since $\phi_{[n,i]}$ are independent, the metric entropy integrability condition on $L_{2}$ is automatically satisfied by Assumption 5. Applying Theorem S1, the random continuous polygon $s_{n}^{\prime}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Phi}_{n}(1))},\frac{\frac{i}{n}\Phi(\frac{i}{n})-\frac{i}{n}\widehat{\Phi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Phi}_{n}(1))}}\right)$
converges to $W$, where $W$ is the standard Wiener process. Then, we have:
$\mathbb{P}\left(W(t)\geq-\gamma^{\ast}(\alpha;k)t^{k},\;\forall\;t\in[0,1]\right)\ =\ 1-\alpha,$ (S5)
where
$\gamma^{\ast}(\alpha;k)=\inf\{\gamma\in\mathbb{R}\mid\mathbb{P}(W(t)\leq\gamma t^{k},\;\forall\;t\in[0,1])\geq 1-\alpha\}.$
Proposition 2 implies that the covariance function $K_{G}$ of the Gaussian Process $G$ satisfies $K_{G}(t,t^{\prime})\geq K_{W}(t,t^{\prime})=0$ where $K_{W}$ is the covariance function of $W$. In addition, we have $K_{G}(t,t)=K_{W}(t,t)$ by construction. Thus, we can apply Slepian’s lemma (Lemma S3) to Equation (S3) and obtain,
$\mathbb{P}\left(G(t)\geq-\gamma^{\ast}(\alpha;k)t^{k},\;\forall\;t\in[0,1]\right)\geq 1-\alpha.$ (S6)
Proposition 1 implies that uniformly for all $p\in[0,1]$, we have as $n\to\infty$:
$\left(\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right)\to(p,G(p))$
Therefore, we can rewrite Equation (S6) as:
$\lim_{n\to\infty}\mathbb{P}\left(\frac{p\Psi(p)-p\widehat{\Psi}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\geq-\gamma^{\ast}(\alpha;k)\left(\frac{\mathbb{V}(p\widehat{\Psi}_{n}(p))}{\mathbb{V}(\widehat{\Psi}_{n}(1))}\right)^{k},\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$
Rearranging the terms gives:
$\lim_{n\to\infty}\mathbb{P}\left(p\Psi(p)\geq p\widehat{\Psi}_{n}(p)-\gamma^{\ast}(\alpha;k)p^{2k}\mathbb{V}(\widehat{\Psi}_{n}(p))^{k}\mathbb{V}(\widehat{\Psi}_{n}(1))^{1/2-k},\;\forall p\in[0,1]\right)\geq 1-\alpha.$
Therefore, for any $c_{0}>0$, we have:
$\lim_{n\to\infty}\mathbb{P}\left(\Psi(p)\geq\widehat{\Psi}_{n}(p)-\gamma(\alpha;k)^{\ast}p^{2k-1}\mathbb{V}(\widehat{\Psi}_{n}(p))^{k}\mathbb{V}(\widehat{\Psi}_{n}(1))^{1/2-k},\;\forall\;p\in[c_{0},1]\right)\geq 1-\alpha.$
Taking $c_{0}\to 0$ gives the required result. ∎
Appendix S5 Proof of Proposition 3
First, we introduce Kahane’s Inequality.
Lemma S4 (Kahane (1985)’s Inequality)
Consider Gaussian random variables $X=(X_{1},\ldots,X_{n})$ and $Y=(Y_{1},\ldots,Y_{n})$ in $\mathbb{R}^{n}$ such that $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]=0$, $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]$, and $\operatorname{Cov}(X_{i},X_{j})\leq\operatorname{Cov}(Y_{i},Y_{j})$ for all $i,j=1,\cdots,n$. Define $Z=(Z_{1},\ldots,Z_{n})$ where $Z_{i}=\sqrt{1-t^{2}}X_{i}+tY_{i}$ for $t\in[0,1]$. If a twice-differentiable function function $f:\mathbb{R}^{n}\to\mathbb{R}$ satisfies for all $i\neq j$ and for all $t\in[0,1]$, i.e., $\mathbb{E}\left[\frac{\partial^{2}f(Z)}{\partial Z_{i}\partial Z_{j}}\right]\geq 0$. Then, we have $\mathbb{E}[f(X)]\leq\mathbb{E}[f(Y)]$.
We then utilize Lemma S4 to prove the following useful “conditional” Slepian inequality.
Lemma S5
For Gaussian random variables $X=(X_{1},\ldots,X_{n})$ and $Y=(Y_{1},\ldots,Y_{n})$ in $\mathbb{R}^{n}$ satisfying $\mathbb{E}[X_{i}]=\mathbb{E}[Y_{i}]=0$, $\mathbb{E}[X_{i}^{2}]=\mathbb{E}[Y_{i}^{2}]$ and $0\leq\operatorname{Cov}(X_{i},X_{j})\leq\operatorname{Cov}(Y_{i},Y_{j})$ for all $i,j=1,\cdots,n$, we have for any $u_{1},\ldots u_{n}\leq 0$:
$\mathbb{P}\left[\bigcap_{i=1}^{n-1}\{X_{i}\geq u_{i}\}\ \Bigl{|}\ X_{n}=0\right]\leq\mathbb{P}\left[\bigcap_{i=1}^{n-1}\{Y_{i}\geq u_{i}\}\ \Bigl{|}\ Y_{n}=0\right].$
Proof.
Consider the function
$f(x_{1},\ldots,x_{n})=\prod_{i=1}^{n-1}\mathbf{1}(x_{i}\geq u_{i})\times\mathbf{1}(x_{n}=0)$
Our goal is to prove $\mathbb{E}[f(X)]\leq\mathbb{E}[f(Y)]$, which immediately implies the desired result. Unfortunately, $f$ is not twice-differentiable, so we approximate it by the following infinitely differentiable function as $\epsilon>0$ approaches 0:
$f_{\epsilon}(x_{1},\ldots,x_{n})=\prod_{i=1}^{n-1}\frac{1}{1+e^{-(x_{i}-u_{i}-\sqrt{\epsilon})/\epsilon}}\times e^{-x_{n}^{2}/\epsilon}$
By construction, we have $\frac{\partial^{2}f_{\epsilon}(x_{1},\ldots,x_{n})}{\partial x_{i}\partial x_{j}}\geq 0$ for all $i,j\neq n$, and so we only need to consider $\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\right]$ where $Z=(Z_{1},\ldots,Z_{n})$ and $Z_{i}=\sqrt{1-t^{2}}X_{i}+tY_{i}$ for $t\in[0,1]$. By definition, we have $\operatorname{Cov}(Z_{i},Z_{j})\geq 0$. Since $u_{i}\leq 0$, we have:
$\mathbb{P}(Z_{i}\leq u_{i},Z_{n}<0)>\mathbb{P}(Z_{i}\leq u_{i},Z_{n}>0)$ (S7)
Now note that the function $f_{\epsilon}(x_{1},\ldots,x_{n})$ is symmetric around zero for $x_{n}$ and approaches the indicator function $\mathbf{1}(x_{i}\geq u_{i})$ from the left for $x_{i}$ satisfying $i\neq n$. Therefore, as $\epsilon$ approaches 0 and $i\neq n$, we have,
$\lim_{\epsilon\downarrow 0}\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\right]$ $=$ $\mathbb{P}(Z_{i}\leq u_{i})\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\ \biggl{|}\ Z_{i}\leq u_{i}\right]$
$=$ $\ \mathbb{P}(Z_{i}\leq u_{i},Z_{n}>0)\underbrace{\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\ \biggl{|}\ Z_{i}\leq u_{i},Z_{n}>0\right]}_{<0}$
$+\mathbb{P}(Z_{i}\leq u_{i},Z_{n}<0)\underbrace{\mathbb{E}\left[\frac{\partial^{2}f_{\epsilon}(Z)}{\partial Z_{i}\partial Z_{n}}\ \biggl{|}\ Z_{i}\leq u_{i},Z_{n}<0\right]}_{>0}$
$>$ $0.$
where the last inequality follows from Equation (S7). Thus, the desired result follows immediately from the application of Lemma S4 to $f_{\epsilon}$ and letting $\epsilon$ apprach 0. ∎
Proposition 1 implies that the random continuous polygon connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right)$
converges to a zero-mean Gaussian process $G(t)$ such that $K_{G}(t,t^{\prime})\geq K_{W}(t,t^{\prime})=0$ where $K_{W}$ is the covariance function of a standard Weiner process $W$. Furthermore, Equation (11) and the continuous mapping theorem imply that the random continuous polygon connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))},\frac{\frac{i}{n}\Psi^{M}(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}}\right)$
converges to a zero-mean Gaussian bridge G(t)tG(1)G(t)-tG(1) on $[0,1]$. By the theory of Gaussian bridges (see e.g., Theorem 3.1 in Sottinen and Yazigi (2014)), we can equivalently characterize this Gaussian bridge as $G(t)\mid G(1)=0$.
Define a set of independent Gaussian variables $\phi_{[n,i]}$ whose mean and variance are identical to those of $\hat{\psi}_{[n,i]}$, i.e., $\mathbb{E}(\phi_{[n,i]})=\mathbb{E}(\hat{\psi}_{[n,i]})$, $\mathbb{V}(\phi_{[n,i]})=\mathbb{V}(\hat{\psi}_{[n,i]})$, and $\operatorname{Cov}(\phi_{[n,i]},\phi_{[n,j]})=0$. Then, define:
$\widehat{\Phi}_{n}(p)$ $=\frac{1}{np}\sum_{i=1}^{\lfloor np\rfloor}\phi_{[n,i]},$
$\widehat{\Phi}^{M}_{n}(p)\$$=\ \widehat{\Phi}_{n}(p)-\frac{p\lfloor np\rfloor}{np}\widehat{\Phi}_{n}(1).$
Theorem S1 implies that the random continuous polygon connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Phi}^{M}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Phi}^{M}_{n}(1))},\frac{\frac{i}{n}\Phi^{M}(\frac{i}{n})-\frac{i}{n}\widehat{\Phi}^{M}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Phi}^{M}_{n}(1))}}\right)$
converges to the Brownian Bridge B(t)=W(t)tW(1)B(t)=W(t)-tW(1). Similarly, this Gaussian process can be characterized as $W(t)\mid W(1)=0$. Then, we have:
$\mathbb{P}\left(W(t)\geq-\delta^{*}_{0}(\alpha)-\delta^{*}_{1}(\alpha)\sqrt{t(1-t)},\;\forall\;t\in[0,1]\mid W(1)=0\right)\ =\ 1-\alpha,$ (S8)
where
$\{\delta^{*}_{0}(\alpha),\delta^{*}_{1}(\alpha)\}$
$=\operatorname*{argmin}_{\delta_{0},\delta_{1}\in\mathbb{R}_{+}^{2}}\left\{\int^{1}_{0}\delta_{0}+\delta_{1}\sqrt{t(1-t)}\ \mathrm{d}t\;\;\text{subject to}\;\;\mathbb{P}(W(t)\leq\delta_{0}+\delta_{1}\sqrt{t(1-t)},\;\forall\;t\in[0,1]\mid W(1)=0)\geq 1-\alpha\right\}.$
The application of Lemma S5 to the stochastic processes $G(t)$ and $W(t)$ implies that we can rewrite Equation (S8) as:
$\mathbb{P}\left(G(t)\geq-\delta^{*}_{0}(\alpha)-\delta^{*}_{1}(\alpha)\sqrt{t(1-t)},\;\forall\;t\in[0,1]\mid G(1)=0\right)\geq 1-\alpha.$ (S9)
Proposition 1 implies that uniformly for all $t\in[0,1]$, we have as $n\to\infty$:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))},\frac{\frac{i}{n}\Psi^{M}(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}^{M}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}}\right)\to(t,G(t)\mid G(1)=0)$
Therefore, we can rewrite Equation (S9) as:
$\lim_{n\to\infty}\mathbb{P}\left(\frac{p\Psi^{M}(p)-p\widehat{\Psi}^{M}_{n}(p)}{\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}}\geq-\delta^{*}_{0}(\alpha)-\delta^{*}_{1}(\alpha)\sqrt{\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)},\;\forall\;p\in[0,1]\right)\geq 1-\alpha.$
Rearranging the terms gives:
$\lim_{n\to\infty}\mathbb{P}\left(p\Psi^{M}(p)\geq p\widehat{\Psi}^{M}_{n}(p)-\delta^{*}_{0}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}-\delta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)},\;\forall p\in[0,1]\right)\geq 1-\alpha.$
Therefore, for any $c_{0}>0$, we have:
$\lim_{n\to\infty}\mathbb{P}\left(\Psi^{M}(p)\geq\widehat{\Psi}^{M}_{n}(p)-\frac{\delta^{*}_{0}(\alpha)}{p}\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}-\delta^{*}_{1}(\alpha)\sqrt{\mathbb{V}(\widehat{\Psi}^{M}_{n}(p))\left(1-\frac{\mathbb{V}(p\widehat{\Psi}^{M}_{n}(p))}{\mathbb{V}(\widehat{\Psi}^{M}_{n}(1))}\right)},\;\forall\;p\in[c_{0},1]\right)\geq 1-\alpha.$
Taking $c_{0}\to 0$ gives the required result. ∎
Appendix S6 Proof of Proposition 4
We begin by proving two lemmas and stating Donsker’s Invariance Principle, which we use to prove Proposition 4.
Lemma S6
Under Assumptions 14A, we have for some $\epsilon>0$:
$\lim_{n\to\infty}n^{1/2+\epsilon}\sup_{p\in[0,1]}\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]\;\mathrm{d}x=0$
$\lim_{n\to\infty}n^{1/2+\epsilon}\sup_{p\in[0,1]}\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid\hat{F}(\hat{S}_{i})=x]\;\mathrm{d}x=0$
Proof.
We first prove the statement for $S_{i}$. By definition of the floor function, we have that $p-\frac{\lfloor np\rfloor}{n}\leq\frac{1}{n}$. Therefore for some $\epsilon<\frac{1}{2}$:
$\lim_{n\to\infty}n^{1/2+\epsilon}\sup_{p\in[0,1]}\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]dx\ \leq\ \lim_{n\to\infty}n^{-1/2+\epsilon}\sup_{p\in[0,1]}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=p]$
By Assumption 4A, we know that $\sup_{p\in[0,1]}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=p]$ is finite. Therefore, the right handside of the above equation is zero, implying the desired result. The same result follows for the statement on $\hat{S}_{i}$. ∎
Lemma S7
Under Assumptions 15, we have
$\sup_{p\in[0,1]}\left(\mathbb{E}\big{[}p\,\widehat{\Psi}_{n}(p)\big{]}-p\,\Psi(p)\right)=o(n^{-1/2}).$
Proof.
Write $\widehat{\Psi}_{n}(p)$ as in (5) with ordering by $\hat{S}_{i}=\hat{f}(\bm{X}_{i})$, and define $A_{p}=\{x:f(x)\geq c(p)\}$, $\widehat{A}_{p}=\{x:\hat{f}(x)\geq\hat{c}(p)\}$, where $c(p)=F^{-1}(1-p)$ and $\hat{c}(p)=\hat{F}^{-1}(1-p)$. By complete randomization (Assumption 2) and the definition of $\psi_{i}$, we have
$\mathbb{E}\big{[}p\,\widehat{\Psi}_{n}(p)\big{]}=\mathbb{E}\!\left[\psi_{i}\,\mathbf{1}\{\bm{X}_{i}\in\widehat{A}_{p}\}\right]+R_{n}(p),\qquad p\,\Psi(p)=\mathbb{E}\!\left[\psi_{i}\,\mathbf{1}\{\bm{X}_{i}\in A_{p}\}\right],$
where $|R_{n}(p)|=o(n^{1/2+\epsilon})$ uniformly in $p$ by Lemma S6 and $\mathbb{E}|\psi_{i}|<\infty$ (Assumption 5). Hence
$\sup_{p\in[0,1]}\Big{|}\,\mathbb{E}\big{[}p\,\widehat{\Psi}_{n}(p)\big{]}-p\,\Psi(p)\,\Big{|}$ $\leq\mathbb{E}|\psi_{i}|\cdot\sup_{p\in[0,1]}\Pr\!\left(\bm{X}_{i}\in A_{p}\triangle\widehat{A}_{p}\right)\;+\;o(n^{-1/2-\epsilon}).$
It suffices to bound $\Pr(\bm{X}\in A_{p}\triangle\widehat{A}_{p})$ uniformly in $p$. Let $r_{n}=\sup_{x}|\hat{f}(x)-f(x)|$. Then we notice that:
$A_{p}\triangle\widehat{A}_{p}\ \subseteq\ \bigl{\{}\,x:\ |f(x)-c(p)|\leq 2r_{n}\,\bigr{\}}.$
Therefore, we have that
$\Pr(\bm{X}\in A_{p}\triangle\widehat{A}_{p})\ \leq\ \Pr\!\left(|f(x)-c(p)|\leq r_{n}\right)\leq Cr_{n}^{\kappa}$
By Assumption 3A, we know that $r_{n}=o(n^{-1/2\kappa})$. Therefore, we have that:
$\sup_{p\in[0,1]}\Pr(\bm{X}\in A_{p}\triangle\widehat{A}_{p})\ =o(n^{-1/2}),$
and the stated bound follows. ∎
Theorem S2 (Donsker’s Invariance Principle)
Consider a double sequence of random variables $U_{i,j}$ with $\mathbb{E}[U_{i,j}]=0$ for all $i,j\in\{1,2,\ldots,n\}$. Assume that $U_{n,i}$ satisfies the metric entropy integrability condition in $L_{2}$ as defined in Equation (3.1) of Ossiander (1987). Then, define the quantity:
$S_{n}(t)=\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}U_{n,i}$
where $t\in[0,1]$ and $S_{n}(0)=0$. Construct the random continuous polygon $s_{n}$ with basic points:
$b_{i}\ =\ \left(\frac{\mathbb{V}(S_{n}(i/n))}{\mathbb{V}(S_{n}(1))},\frac{S_{n}(i/n)}{\sqrt{\mathbb{V}(S_{n}(1))}}\right),$
where $i\in\{1,\cdots,n\}$. Then, as the sample size $n$ goes to infinity, $s_{n}$ converges in distribution to $G$ uniformly in the Skorokhod space, where $G$ is a zero-mean Gaussian process with covariance function
$K(p,p^{\prime})=\lim_{n\to\infty}\operatorname{Cov}\left(\frac{S_{n}(\lfloor pn\rfloor/n)}{\sqrt{\mathbb{V}(S_{n}(\lfloor pn\rfloor/n))}},\frac{S_{n}(\lfloor p^{\prime}n\rfloor/n)}{\sqrt{\mathbb{V}(S_{n}(\lfloor p^{\prime}n\rfloor/n))}}\right),$
in particular $K(p,p)=p$. Additionally, when $\operatorname{Cov}(U_{n,i},U_{n,j})=0$ for all $n$ and $1\leq i,j\leq n$, then $G$ is the standard Wiener process.
Finally, using the above results, we prove Proposition 4. Define for all $i\in\{1,\cdots,n\}$:
$\Delta\Psi_{n,i}=\frac{i}{n}\mathbb{E}\left[\psi_{i}\ \biggl{|}\ F(S_{i})\geq\frac{i}{n}\right]-\frac{i-1}{n}\mathbb{E}\left[\psi_{i}\ \biggl{|}\ F(S_{i})\geq\frac{i-1}{n}\right].$
Note that:
$p\widehat{\Psi}_{n}(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\hat{\psi}_{[n,i]},\qquad p\Psi(p)=\frac{1}{n}\sum_{i=1}^{\lfloor np\rfloor}\Delta\Psi_{n,i}+\int^{1-p}_{\frac{\lfloor n(1-p)\rfloor}{n}}\mathbb{E}[\hat{\psi}_{i}\mid F(S_{i})=x]\;\mathrm{d}x$
By Lemmas S7 and S6, we can construct a new random variable $U_{n,i}$ with $\mathbb{E}(U_{n,i})=0$ for some fixed quantity $\epsilon_{n,i}>0$ in the following manner,
$U_{n,i}$ $=\Delta\Psi_{n,i}-\hat{\psi}_{[n,i]}-\epsilon_{n,i},\quad\text{where}\ \sup_{t\in[0,1]}\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\epsilon_{n,i}=o(n^{-1/2}).$ (S10)
By Assumption 5 $\psi_{n,i}$ have finite second moments, so the induced order statistics $\psi_{[n,i]}$ satisfy the metric entropy integrability condition on $L_{2}$ (Davydov and Egorov, 2000). Therefore we can apply Theorem S1 to $U_{n,i}$ by considering a random continuous polygon $\tilde{s}_{n}$ connected by the basic points:
$\tilde{b}_{i}\ =\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\hat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})-\frac{1}{n}\sum_{i^{\prime}=1}^{i}\epsilon_{n,i^{\prime}}}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right),$
implying that $\tilde{s}_{n}$ converges to a zero-mean Gaussian Process. Equation (S10) implies $\frac{\frac{1}{n}\sum_{i=1}^{\lfloor nt\rfloor}\epsilon_{n,i}}{\sqrt{\mathbb{V}(\Psi(1))}}\to 0$ uniformly in $t$. Thus, the random continuous polygon $s_{n}$ connected by the basic points:
$\left(\frac{\mathbb{V}(\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n}))}{\mathbb{V}(\widehat{\Psi}_{n}(1))},\frac{\frac{i}{n}\Psi(\frac{i}{n})-\frac{i}{n}\widehat{\Psi}_{n}(\frac{i}{n})}{\sqrt{\mathbb{V}(\widehat{\Psi}_{n}(1))}}\right),$
converges to a zero-mean Gaussian process $G$. ∎